### v2

# Iterated sampling importance resampling with adaptive number of proposals

Abstract.
Iterated sampling importance resampling (i-SIR) is a Markov chain Monte Carlo (MCMC) algorithm which is based on $N$ independent proposals. As $N$ grows, its samples become nearly independent, but with an increased computational cost. We discuss a method which finds an approximately optimal number of proposals $N$ in terms of the asymptotic efficiency. The optimal $N$ depends on both the mixing properties of the i-SIR chain and the (parallel) computing costs. Our method for finding an appropriate $N$ is based on an approximate asymptotic variance of the i-SIR, which has similar properties as the i-SIR asymptotic variance, and a generalised i-SIR transition having fractional ‘number of proposals.’ These lead to an adaptive i-SIR algorithm, which tunes the number of proposals automatically during sampling. Our experiments demonstrate that our approximate efficiency and the adaptive i-SIR algorithm have promising empirical behaviour. We also present new theoretical results regarding the i-SIR, such as the convexity of asymptotic variance in the number of proposals, which can be of independent interest.
1. Introduction
Importance sampling is an ubiquitous method in Bayesian statistics which can be used to estimate expectations of the form
$\pi(f)=\mathbb{E}_{\pi}[f(X)]=\int_{\mathbb{X}}f(x)\pi(x)\mathrm{d}x,$
where $\pi$ is a probability density on a state space $\mathbb{X}$, [^1] Usually, $\mathbb{X}\subset\mathbb{R}^{d}$ and the integral is with respect to the Lebesgue measure, but the method works with any underlying space $\mathbb{X}$ equipped with a $\sigma$-finite dominating measure $\mathrm{d}x$’. and $f:\mathbb{X}\to\mathbb{R}$ is a (measurable) test function. In importance sampling, $N$ independent samples are drawn from a ‘proposal’ distribution $q$ which dominates $\pi$, [^2] That is, $\pi$ is absolutely continuous with respect to $q$, or $q(x)=0$ implies $\pi(x)=0$. and the samples $X_{i}$ are associated with importance weights $W_{i}\propto\pi(X_{i})/q(X_{i})$.
Iterated sampling importance resampling (i-SIR) algorithm, which was proposed in [tjelmeland] and generalised to a conditional sequential Monte Carlo algorithm in [andrieu-doucet-holenstein], is a Markov chain Monte Carlo (MCMC) method, which uses the same ingredients as importance sampling, but is iterated $n$ times with fixed (moderate) $N$, and leads to an (approximate and correlated) sample $X_{1},\ldots,X_{n}\sim\pi$.
In contrast with importance sampling, the i-SIR produces an unweighted sample, which can be more convenient in some applications. The i-SIR can also be used as ‘burn-in’ to (self-normalised) importance sampling in order to mitigate its bias; this aspect was investigated in [cardoso-samsonov-thin-moulines-olsson]. The i-SIR can also be readily used within a component-wise MCMC sampler; for instance, in a latent variable model, i-SIR can serve as a Markov update for latent variables given hyperparameters. In this context, there is another strategy to use importance sampling within MCMC, known as the ‘grouped independent Metropolis–Hastings’ (GIMH) [andrieu-roberts]. It is known to suffer from problematic mixing behaviour, which does not improve as $N$ increases, contrary to the i-SIR [andrieu-vihola-pseudo, andrieu-lee-vihola].
We present an adaptive i-SIR algorithm, which automatically selects the number of proposals $N$, which minimises the cost weighted approximate asymptotic variance [glynn-whitt]. We start by presenting the i-SIR and its continuous-number-of-proposals relaxation (Section 2) and summarise some of its theoretical properties (Section 3). In particular, Theorem 4 guarantees that the asymptotic variance of i-SIR is strictly decreasing and convex with respect to the number of proposals. We then discuss its efficiency and present an approximate efficiency in Section 4, which our adaptive algorithm discussed in Section 5 aims to minimise. We prove a strong law of large numbers for our adaptive i-SIR algorithm (Theorem 7), which guarantees consistency of estimates based on its output.
Sections 69 are devoted to technical results required by our theoretical results. Section 6 is about the behaviour of stationary covariances and acceptance rates, which lead to the convexity of the asymptotic variance in Section 7. Section 8 presents upper and lower bounds for i-SIR acceptance rate and asymptotic variance, and Section 9 discusses i-SIR in discrete state spaces, and its alternative approximation.
Our empirical findings in Section 10 suggest that our approximate efficiency criterion can be quite robust, but also exposes some problematic cases. Our adaptive algorithm shows promising behaviour in our experiments. Section 11 concludes with a discussion, including possible directions in which our algorithm could be extended.
2. The i-SIR algorithm
The i-SIR algorithm defines a Markov chain Monte Carlo method targeting a probability density $\pi(x)$. It uses samples from a proposal probability density $q(x)$, which is assumed to dominate $\pi(x)$, that is, $q(x)=0$ implies $\pi(x)=0$. Then, the (unnormalised) importance weights $w(x)=c\pi(x)/q(x)\in[0,\infty)$ are well-defined for all $x$, where we use the convention $0/0=0$. The value of the constant $c>0$ need not be known; typically it stems from that we can evaluate only the unnormalised target density $\pi_{u}(x)=c\pi(x)$, in which case $c=(\int\pi_{u}(x)\mathrm{d}x)^{-1}$ is the unknown normalising constant.
Algorithm 1 summarises the iteration $k$ of i-SIR, that is, the Markov transition $X_{k-1}\to X_{k}$. In Algorithm 1 and hereafter, we use the notation i:ji{:}j to denote a range of integers, that is, $i{:}j=(i,i+1,\ldots,j-1,j)$ to mean integers from $i$ to $j$ and $I\sim\mathrm{Categorical}(w^{1:N})$ means a categorical distribution with weights $w^{1},\ldots,w^{N}$, that is, $\mathbb{P}(I=k)=w^{k}$.
Algorithm 1 $\textsc{iSIR}(X_{k-1},N)$
1:Draw independent $Y_{k}^{2:N}\sim q$ and set $Y_{k}^{1}=X_{k-1}$
2:Calculate weights $W_{k}^{i}=w(X_{k}^{i})/\sum_{j=1}^{N}w(X_{k}^{j})$
3:Sample $I\sim\mathrm{Categorical}(W_{k}^{1:N})$
4:Return $X_{k}=Y_{k}^{I}$
The i-SIR algorithm does not involve an explicit ‘accept/reject’ decision; instead, sampling $I=1$ leads to a ‘rejection’ because $X_{k}=X_{k-1}$. In the case $N=2$, the i-SIR coincides with independence sampler having Barker’s acceptance rate [barker]. We focus on the i-SIR of Algorithm 1, but note that the algorithm can be extended to cover dependent proposals [e.g. mendes-scharth-kohn, karppinen-vihola] with similar flavour as some other multiple-try Metropolis algorithms [see, e.g., the review martino-review].
It is not hard to check that the i-SIR algorithm has the following transition probability:
(1) $P_{N}(z_{1},A)=\int\bigg(\prod_{n=2}^{N}q(z_{n})\bigg)\sum_{i=1}^{N}\frac{w(z_{i})}{\sum_{j=1}^{N}w(z_{j})}\mathbf{1}\left(z_{i}\in A\right)\mathrm{d}z_{2}\cdots\mathrm{d}z_{n}.$
We present a simple ‘continuous relaxation’ of the i-SIR algorithm, when we discuss the adaptation of the number of proposals in Section 5. It is defined for a ‘fractional number of proposals’ $\lambda\geq 1$ (although later we will focus on $\lambda\geq 2$ only). Namely, let $\lambda\geq 1$ and $N$ be an integer such that $N\leq\lambda<N+1$, then
$P_{\lambda}(x,A)=\beta P_{N}(x,A)+(1-\beta)P_{N+1}(x,A),$
where the interpolation constant $\beta=N+1-\lambda\in(0,1]$. Algorithm 2 presents a particular way of sampling from $P_{\lambda}$, which will be useful later.
Algorithm 2 $\textsc{iSIR}(X_{k-1},\lambda)$
1:Draw independent $Y_{k}^{2:\bar{N}_{k}}\sim q$ where $\bar{N}_{k}=\lfloor\lambda\rfloor+1$
2:With probability $\beta_{k}=\bar{N}_{k}-\lambda$ set $N_{k}=\bar{N}_{k}-1$, else set $N_{k}=\bar{N}_{k}$, and calculate weights $W_{k}^{i}=w(X_{k}^{i})/\sum_{j=1}^{N_{k}}w(X_{k}^{j})$
3:Sample $I\sim\mathrm{Categorical}(W_{k}^{1:N_{k}})$
4:Return $X_{k}=Y_{k}^{I}$
3. Properties of the i-SIR
It is well-known that the i-SIR algorithm $P_{N}$ is reversible with respect to $\pi$ [e.g. tjelmeland, andrieu-lee-vihola], and therefore the generalised i-SIR $P_{\lambda}$, as a convex combination of $P_{N}$ and $P_{N+1}$ is $\pi$-reversible. We assume the following to hold throughout Sections 35:
Assumption (A1).
The weight function is bounded:
$\hat{w}=\sup_{x\in\mathbb{X}}w(x)<\infty.$
This assumption requires, roughly speaking, the target density $\pi$ to have lighter tails than the proposal $q$, which can often be guaranteed by a suitable choice of $q$. It is known that i-SIR is slowly mixing (subgeometric) essentially if (A1) does not hold, and if (A1) holds, then the i-SIR is guaranteed to be well-behaved (uniformly ergodic) [cf. andrieu-lee-vihola]. In fact, for large $\lambda$, $P_{\lambda}$ corresponds asymptotically to independent sampling from $\pi$. Hereafter, we denote by $\pi$ the probability measure corresponding to the density $\pi$, that is, $\pi(A)=\int_{A}\pi(x)\mathrm{d}x$.
Proposition 1.
The i-SIR transition probability $P_{\lambda}$ satisfies the following strong minorisation condition:
$P_{\lambda}(x,A)\geq\frac{\lambda-1}{2\hat{w}+\lambda-1}\pi(A).$
This result is from [andrieu-lee-vihola] for integer $\lambda$ and detailed in Lemma 49 in Appendix E for non-integer $\lambda$.
Our main focus will be on the asymptotic variance of the i-SIR chain, defined for finite-variance functionals. Hereafter, we denote the integral of a (measurable) function $f:\mathbb{X}\to\mathbb{R}$ with respect to $\pi$ by $\pi(f)=\int f(x)\pi(x)\mathrm{d}x$ and by $L^{2}(\pi)$ the set of all finite variance functionals with respect to $\pi$, that is, such that $\pi(f^{2})<\infty$, and also define $\mathrm{var}_{\pi}(f)=\pi(f^{2})-\pi(f)^{2}$.
Definition 2.
Fix $\lambda\in(1,\infty)$ and assume that $(X_{k})_{k\geq 0}$ is the stationary i-SIR chain following $P_{\lambda}$, that is, $X_{0}\sim\pi$ and $X_{k}\mid(X_{k-1}=x)\sim P_{\lambda}(x,\;\cdot\;)$. The asymptotic variance of $f\in L^{2}(\pi)$ is:
$\mathrm{var}(P_{\lambda},f)=\mathrm{var}\big(f(X_{0})\big)+2\sum_{k=1}^{\infty}\mathrm{cov}\big(f(X_{0}),f(X_{k})\big).$
Thanks to the uniform ergodicity of $P_{\lambda}$ ensured by Proposition 1, $\mathrm{var}(P_{\lambda},f)$ is always well-defined and finite. Moreover, a central limit theorem holds [e.g. jones-survey, Theorem 9]:
Theorem 3.
For $\lambda\in(1,\infty)$, let $X_{k}$ be the i-SIR chain with an arbitrary initial state $X_{0}=x$. Then, for any $f\in L^{2}(\pi)$:
$\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\big[f(X_{k})-\pi(f)\big]\xrightarrow{n\to\infty}N\big(0,\mathrm{var}(P_{\lambda},f)\big).$
Theorem 3 shows that $\mathrm{var}(P_{\lambda},f)$ quantifies the statistical (per sample) efficiency of estimating $\pi(f)$ with a Markov chain following $P_{\lambda}$. While we do not know $\mathrm{var}(P_{\lambda},f)$ in general, we know that it has certain properties.
Theorem 4.
Let $\lambda\in(1,\infty)$ and $f\in L^{2}(\pi)$ such that $\mathrm{var}_{\pi}(f)>0$. Then,
  1. (i)
    $\mathrm{var}_{\pi}(f)\leq\mathrm{var}(P_{\lambda},f)\leq\frac{4\hat{w}+\lambda-1}{\lambda-1}\mathrm{var}_{\pi}(f)$.
  2. (ii)
    $\lambda\mapsto\mathrm{var}(P_{\lambda},f)$ is strictly convex and decreasing on $(1,\infty)$.
Theorem 4 (i) is similar to the bound in [andrieu-lee-vihola], and follows from Proposition 1; see Lemma 33.Theorem 4 (ii) is to our knowledge new and given in Theorem 27.
4. Overall efficiency and approximation based on average acceptance rate
The asymptotic variance measures the statistical efficiency of $P_{\lambda}$, but because sampling $P_{\lambda}$ involves simulation of $O(\lambda)$ samples and evaluation of the weight function at each simulated sample, the number of computer operations is $O(\lambda)$. This affects the overall efficiency of $P_{\lambda}$. However, thanks to the independence of the samples, the sampling and evaluation of the weight can be performed in parallel, so number of computer operations might not be directly useful.
We will assume that $c(\lambda)$ is a function that defines the ‘cost’ of (time spent on) one iteration of $P_{\lambda}$. In particular, we will focus on a simplified cost of the form $c(\lambda)=a+b\lambda$, which assumes an initial (‘parallelisation overhead’) cost $a>0$ and a term which increases in terms of $\lambda$ (‘parallel computing cost’), with rate $b>0$; see [gustafson] for other ways to characterise parallel computing speedups.
When we take into account the $\lambda$-dependent cost $c(\lambda)$, we would ultimately want to minimise the so-called inverse relative efficiency $L(\lambda)=c(\lambda)\mathrm{var}(P_{\lambda},f)$ in terms of a test function $f$ of interest [cf. glynn-whitt]. However, this is difficult in practice, for two reasons. The asymptotic variance $\mathrm{var}(P_{\lambda},f)$ is unknown, and notoriously difficult to estimate. Furthermore, we often have multiple test functions $f_{1},\ldots,f_{m}$ which we want to estimate, which will generally have different asymptotic variances.
The upper bound of the asymptotic variance (i) might be a possible proxy to optimise. However, while the existence of $\hat{w}$$\infty$ in (A1) might be known, its value is rarely known. It is also unclear how well the upper bound matches the actual asymptotic variance. (We present some numerical comparisons in Section 10 regarding this.)
Instead of the asymptotic variance upper bound, we consider the following approximation of the i-SIR transition inspired by Proposition 1:
Definition 5.
The approximate i-SIR transition $\tilde{P}_{\lambda}$ of $P_{\lambda}$ is defined as
$\tilde{P}_{\lambda}(x,A)=\big(1-\epsilon(\lambda)\big)\pi(A)+\epsilon(\lambda)\delta_{x}(A),$
where the average holding probability $\epsilon(\lambda)$ is
$\epsilon(\lambda)=\int\pi(x)P_{\lambda}(x,\{x\})\mathrm{d}x.$
In words, $\tilde{P}_{\lambda}$ corresponds to independent sampling with probability $1-\epsilon(\lambda)$ and ‘rejection’ (not moving) with probability $\epsilon(\lambda)$. Proposition 1 guarantees that we can write $P_{\lambda}(x,A)=(1-\delta(\lambda))\pi(A)+\delta(\lambda)Q_{\lambda}(x,A)$, where $\delta(\lambda)\to 0$ as $\lambda\to\infty$, and the ‘residual’ kernel $Q_{\lambda}(x,A)$ is $\pi$-reversible. This suggests that at least for $\lambda$ large, $\tilde{P}_{\lambda}\approx P_{\lambda}$.
The approximate transition probability $\tilde{P}_{\lambda}$ is simple enough to admit an explicit form of asymptotic variance:
Proposition 6.
For any $f\in L^{2}(\pi)$ such that $\mathrm{var}_{\pi}(f)>0$
  1. (i)
    $\mathrm{var}(\tilde{P}_{\lambda},f)=\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)}\mathrm{var}_{\pi}(f).$
  2. (ii)
    $\mathrm{var}_{\pi}(f)\leq\mathrm{var}(\tilde{P}_{\lambda},f)\leq\frac{4\hat{w}+\lambda-1}{\lambda-1}\mathrm{var}_{\pi}(f)$
  3. (iii)
    $\lambda\mapsto\mathrm{var}(\tilde{P}_{\lambda},f)$ is strictly convex and decreasing.
Proposition 6 (i) follows from Lemma 37 in Appendix B, (ii) from Lemma 29 and (iii) from Lemma 24. Proposition 6 indicates that $\mathrm{var}(\tilde{P}_{\lambda},f)$ has similar properties as $\mathrm{var}(P_{\lambda},f)$ (cf. Theorem 4). Our method for finding an appropriate ‘number of proposals’ $\lambda$ is based on this expression: we aim for approximately minimising $c(\lambda)\mathrm{var}(\tilde{P}_{\lambda},f)$, which we discuss next.
5. Adaptation of the number of proposals
Our method for finding an appropriate $\lambda$ is a stochastic gradient type algorithm. To simplify discussion in this section, we focus on a continuous case (e.g. $\mathbb{X}\subset\mathbb{R}^{d}$), and assume following to hold:
Assumption (A2).
The proposal distribution $q$ is non-atomic, that is, if $Y\sim q$, then $\mathbb{P}(Y=c)=0$ for all $c\in\mathbb{X}$.
The i-SIR is valid also when $q$ has atoms, like when $\mathbb{X}$ is discrete; we include discussion about this case in Sections 69.
Under (A2), the i-SIR chain has $X_{k}=X_{k-1}$ exactly when $I=1$, and therefore, the average holding (or rejection) probability of the chain is:
$\epsilon(\lambda)=\int\pi(z_{1})\bigg(\prod_{i=2}^{N(\lambda)}q(z_{i})\bigg)w(z_{1})\bigg(\frac{\beta(\lambda)}{\sum_{j=1}^{N(\lambda)-1}w(z_{j})}+\frac{1-\beta(\lambda)}{\sum_{j=1}^{N(\lambda)}w(z_{j})}\bigg)\mathrm{d}z_{1}\cdots\mathrm{d}z_{N(\lambda)},$
where $\beta(\lambda)=\lfloor\lambda\rfloor+1-\lambda$ and $N(\lambda)=\lfloor\lambda\rfloor+1$ Clearly, the function $\epsilon$ is differentiable at non-integer points $\lambda$:
(2) $\epsilon^{\prime}(\lambda)$ $=\int\pi(z_{1})\bigg(\prod_{i=2}^{N(\lambda)}q(z_{i})\bigg)w(z_{1})\bigg(\frac{1}{\sum_{j=1}^{N(\lambda)}w(z_{j})}-\frac{1}{\sum_{j=1}^{N(\lambda)-1}w(z_{j})}\bigg)\mathrm{d}z_{1}\cdots\mathrm{d}z_{N(\lambda)},$
and we extend $\epsilon^{\prime}$ to all $\lambda\geq 1$ by taking the cádlág extension of the right-hand side of (2). If we iterated Algorithm 2 (with fixed $\lambda$), we could estimate $\epsilon(\lambda)$ and $\epsilon^{\prime}(\lambda)$ by averaging the following quantities:
(3) $\hat{\epsilon}_{k}(\lambda)$ $=w(Y_{k}^{1})\bigg(\frac{\beta_{k}}{\sum_{j=1}^{\bar{N}_{k}-1}w(Y_{k}^{j})}+\frac{1-\beta_{k}}{\sum_{j=1}^{\bar{N}_{k}}w(Y_{k}^{j})}\bigg)$
(4) $\hat{\epsilon}^{\prime}_{k}(\lambda)$ $=w(Y_{k}^{1})\bigg(\frac{1}{\sum_{j=1}^{\bar{N}_{k}}w(Y_{k}^{j})}-\frac{1}{\sum_{j=1}^{\bar{N}_{k}-1}w(Y_{k}^{j})}\bigg).$
Under stationarity, that is, when $X_{k-1}\sim\pi$, it is easy to see that the expected values of $\hat{\epsilon}_{k}(\lambda)$ and $\hat{\epsilon}^{\prime}_{k}(\lambda)$ are indeed $\epsilon(\lambda)$ and $\epsilon^{\prime}(\lambda)$, respectively. Ignoring the multiplicative constant $\mathrm{var}_{\pi}(f)$ in $\mathrm{var}(\tilde{P}_{\lambda},f)$, which does not depend on $\lambda$, we consider the following loss function:
(5) $\tilde{L}(\lambda)=\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)}c(\lambda),$
which is continuous and differentiable (at non-integer $\lambda$):
$\tilde{L}^{\prime}(\lambda)$ $=c^{\prime}(\lambda)\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)}+c(\lambda)\frac{\epsilon^{\prime}(\lambda)\big(1-\epsilon(\lambda)\big)+\epsilon^{\prime}(\lambda)\big(1+\epsilon(\lambda)\big)}{\big(1-\epsilon(\lambda)\big)^{2}}$
$=\frac{c^{\prime}(\lambda)(1-\epsilon^{2}(\lambda))+2c(\lambda)\epsilon^{\prime}(\lambda)}{\big(1-\epsilon(\lambda)\big)^{2}}.$
For a gradient-like minimisation algorithm, we may ignore the positive denominator, which scales the gradient. Furthermore, we simplify $\epsilon^{2}(\lambda)\approx\epsilon_{s}(\lambda)$, which is where ergodic averages of the squares $\hat{\epsilon}^{2}_{k}(\lambda)$ of (3) converge; see (12) in Section 8. Thanks to Jensen’s inequality, $\epsilon^{2}(\lambda)\leq\epsilon_{s}(\lambda)$.
We end up with a rescaled approximate gradient of the form
$G(\lambda)=c^{\prime}(\lambda)(1-\epsilon_{s}(\lambda))+2c(\lambda)\epsilon^{\prime}(\lambda).$
Furthermore, when we adapt $\lambda$, we use a monotone reparametrisation $\lambda=1+e^{\xi}$, where $\xi\in\mathbb{R}$ is the parameter we adapt.
One iteration of the algorithm, that is, update $(\xi_{k-1},X_{k-1})$ to $(\xi_{k},X_{k})$ is summarised in Algorithm 3.
Algorithm 3 $\textsc{iSIRAdapt}(X_{k-1},\xi_{k-1})$
1:Let $\lambda_{k-1}=e^{\xi_{k-1}}+1$
2:Let $Y_{k}^{1}=X_{k-1}$ and simulate $Y_{k}^{2:(\lfloor\lambda_{k-1}\rfloor+1)}\overset{\text{i.i.d.}}{\sim}q$
3:With probability $\lambda_{k-1}-\lfloor\lambda_{k-1}\rfloor$ set $N_{k}=\lfloor\lambda_{k-1}\rfloor+1$, else $N_{k}=\lfloor\lambda_{k-1}\rfloor$
4:Draw $I\sim\mathrm{Categ}(w(Y_{k}^{1}),\ldots,w(Y_{k}^{N_{k}}))$
5:Set $X_{k}=Y_{k}^{I}$
6:Calculate $\hat{\epsilon}_{k}(\lambda_{k-1})$ and $\hat{\epsilon}^{\prime}_{k}(\lambda_{k-1})$ as in (3) and (4), respectively
7:Calculate $\xi_{k}=\mathrm{Proj}\big\{\xi_{k-1}-\gamma_{k}\big[c^{\prime}(\lambda_{k-1})\big(1-\hat{\epsilon}_{k}(\lambda_{k-1})^{2}\big)+2c(\lambda_{k-1})\hat{\epsilon}^{\prime}_{k}(\lambda_{k-1})\big]\big\}$
The algorithm uses stochastic observations, whose stationary expectations coincide with $G$. It is in instance of stochastic approximation type adaptive MCMC [cf. andrieu-robert, andrieu-moulines, andrieu-thoms] where $\gamma_{k}$ is a constant step size sequence, which we choose as $k^{-\beta}$ where $\beta=0.75$; values $\beta\in(1/2,1)$ are advisable [e.g. andrieu-moulines]. The additional projection function $\mathrm{Proj}\{\xi\}=\min\big\{\max\{0,\xi\},\log(N_{\max}-1)\big\}$ ensures that $2\leq\lambda_{k}\leq N_{\max}$ when $N_{\max}$ is finite, and $2\leq\lambda_{k}<\infty$ when $N_{\max}=\infty$. A finite upper bound is not necessary for the validity, but it can simplify the practical implementation of the method. The initial value of $\xi_{0}$ is arbitrary, and can correspond for instance to $\lambda_{0}=N_{\max}/2$.
Theorem 7.
Suppose that $\gamma_{k}=k^{-\beta}$ where $\beta>0$. Suppose the cost function $c$ is continuous and differentiable, and in the case $N_{\max}=\infty$, additionally $c(\lambda)=O(\lambda)$ and $c^{\prime}$ is bounded. Then, the variables $X_{1},X_{2},\ldots$ generated by iteration of Algorithm 3 satisfy the following strong law of large numbers for any bounded (measurable) test function $f$:
$\frac{1}{n}\sum_{k=1}^{n}f(X_{k})\xrightarrow{n\to\infty}\pi(f).$
Proof.
When $N_{\max}$ is finite, we set $A=[0,\log(N_{\max}-1)]$, and when $N_{\max}=\infty$, we set $A=[0,\infty)$. Let us denote
$H_{k}=-\big[c^{\prime}(\lambda_{k-1})\big(1-\hat{\epsilon}_{k}(\lambda_{k-1})^{2}\big)+2c(\lambda_{k-1})\hat{\epsilon}^{\prime}_{k}(\lambda_{k-1})\big].$
By definition
$\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})=\begin{cases}\xi_{k-1}+\gamma_{k}H_{k},&\xi_{k-1}+\gamma_{k}H_{k}\in A,\\ 0,&\xi_{k-1}+\gamma_{k}H_{k}<0\\ \log(N_{\max}-1),&\xi_{k-1}+\gamma_{k}H_{k}>\log(N_{\max}-1).\end{cases}$
Let $\tilde{H}_{k}=\frac{\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})-\xi_{k-1}}{\gamma_{k}}$, then we obtain
$\xi_{k}=\mathrm{Proj}(\xi_{k-1}+\gamma_{k}H_{k})=\xi_{k-1}+\gamma_{k}\tilde{H}_{k},$
$|\xi_{k}-\xi_{k-1}|=\gamma_{k}|\tilde{H}_{k}|$ and clearly $\gamma_{k}|\tilde{H}_{k}|\leq\gamma_{k}|H_{k}|$. We have $|1-\hat{\epsilon}_{k}(\lambda_{k-1})^{2}|\leq 1$. Now $\lambda_{k-1}=e^{\xi_{k-1}}+1$, let us denote $N(x)=\lfloor e^{x}+1\rfloor$, then
$\mathbb{E}\big[|\hat{\epsilon}_{k}^{\prime}(\lambda_{k-1})|\;\big|\;X_{k-1},\xi_{k-1}\big]$ $=\mathbb{E}\Bigg[\bigg|\frac{w(Y_{k}^{1})}{\sum_{j=1}^{N(\xi_{k-1})+1}w(Y_{k}^{j})}-\frac{w(Y_{k}^{1})}{\sum_{j=1}^{N(\xi_{k-1})}w(Y_{k}^{j})}\bigg|\;\Bigg|\;X_{k-1},\xi_{k-1}\Bigg]$
$=\mathbb{E}\bigg[\frac{w(Y_{k}^{1})w(Y_{k}^{N(\xi_{k-1})+1})}{\big(\sum_{i=1}^{N(\xi_{k-1})}w(Y_{k}^{i})\big)\big(\sum_{j=1}^{N(\xi_{k-1})+1}w(Y_{k}^{j})\big)}\;\bigg|\;X_{k-1},\xi_{k-1}\bigg]$
$\leq\mathbb{E}\bigg[\frac{w(Y_{k}^{N(\xi_{k-1})+1})}{\sum_{j=1}^{N(\xi_{k-1})+1}w(Y_{k}^{j})}\;\bigg|\;X_{k-1},\xi_{k-1}\bigg]$
$=\mathbb{E}\bigg[\frac{1}{N(\xi_{k-1})}\frac{\sum_{i=2}^{N(\xi_{k-1})+1}w(Y_{k}^{i})}{\sum_{j=1}^{N(\xi_{k-1})+1}w(Y_{k}^{j})}\;\bigg|\;X_{k-1},\xi_{k-1}\bigg]\leq\frac{1}{N(\xi_{k-1})},$
where the last equality follows from the fact that $Y_{2},\dots,Y_{N}$ are exchangeable for fixed $N$. Therefore
$\mathbb{E}\big[|c(\lambda_{k-1})\hat{\epsilon}_{k}^{\prime}(\lambda_{k-1})|\;\big|\;X_{k-1},\xi_{k-1}\big]\leq\frac{|c(e^{\xi_{k-1}}+1)|}{\lfloor e^{\xi_{k-1}}+1\rfloor}\leq\frac{|c(e^{\xi_{k-1}}+1)|}{e^{\xi_{k-1}}}.$
When $N_{\max}$ is finite, functions $x\mapsto\frac{|c(e^{x}+1)|}{e^{x}}$ and $x\mapsto c^{\prime}(e^{x})$ are bounded on $A$. When $N_{\max}=\infty$ and the additional assumptions hold, then $x\mapsto\frac{|c(e^{x}+1)|}{e^{x}}$ is continuous and $\frac{|c(e^{x}+1)|}{e^{x}}=O(1)$, hence bounded on $A$. Thus, in both cases
$\mathbb{E}\big[|H_{k}|\big]\leq\mathbb{E}\big[|c^{\prime}(\lambda_{k-1})|(1-\hat{\epsilon}_{k}(\lambda_{k-1})^{2}\big]+2\mathbb{E}\big[\mathbb{E}\big[|c(\lambda_{k-1})\hat{\epsilon}^{\prime}_{k}(\lambda_{k-1})|\;\big|\;X_{k-1},\xi_{k-1}\big]\big]\leq M,$
for some $M>0$, hence
$\sum_{k=1}^{\infty}\frac{\gamma_{k}}{k}\mathbb{E}\big[|\tilde{H}_{k}|\big]\leq\sum_{k=1}^{\infty}\frac{\gamma_{k}}{k}\mathbb{E}\big[|H_{k}|\big]=\sum_{k=1}^{\infty}k^{-(\beta+1)}\mathbb{E}\big[|H_{k}|\big]\leq M\sum_{k=1}^{\infty}k^{-(\beta+1)}<\infty.$
Let $\xi,\xi^{\prime}\in A$, then by Lemma 49
$d_{\mathrm{tv}}\big(P_{e^{\xi^{\prime}}+1}(x,\;\cdot\;),P_{e^{\xi}+1}(x,\;\cdot\;)\big)$ $\leq\frac{6|e^{\xi^{\prime}}-e^{\xi}|}{\max\{e^{\xi^{\prime}}+1,e^{\xi}+1\}}\leq 6|1-e^{-|\xi^{\prime}-\xi|}|\leq 6|\xi^{\prime}-\xi|,$
where the last inequality follows since $x\mapsto x-(1-e^{-x})$ is non-negative and increasing on $0,\infty)$. Therefore, by combining all the previous and Lemma 50 in Appendix E with [laitinen-vihola-adaptive, Lemma 16 $(ii)$, Theorem 9 $(ii)$], we obtain strong law of large numbers. ∎
Theorem 7 ensures that the simulated samples lead to consistent inference, but does not guarantee the convergence of the adaptation, that is, whether there exists a limit $\lambda_{*}$ such that $\lambda_{k}\to\lambda_{*}$. Our empirical observations in Section 10 suggest that such convergence happens, and that $\lambda_{*}$ does indeed (approximately) minimise the loss in (5) (and also approximately minimises the inverse relative efficiency). However, precise theoretical results about this convergence are out of the scope of this paper. We conclude this section with a discussion of some partial results related to the convergence analysis.
Here, we consider a gradient descent like stochastic approximation for an unbounded interval $\lambda\in(1,\infty)$. If it converges, the limit point is the zero of the ‘gradient-like’ function $G(\lambda)$.
Proposition 8.
Let $c(\lambda)=a+b\lambda$, $a\geq 0$, $b>0$ and $h_{1},h_{2},h_{3}:[1,\infty)\to\mathbb{R}$,
$h_{1}(\lambda)$ $=c^{\prime}(\lambda)(1-\lambda^{-2})+2c(\lambda)\epsilon^{\prime}(\lambda)$
$h_{2}(\lambda)$ $=c^{\prime}(\lambda)(1-\epsilon(\lambda)^{2})+2c(\lambda)\epsilon^{\prime}(\lambda)$
$h_{3}(\lambda)$ $=c^{\prime}(\lambda)(1-\epsilon_{s}(\lambda))+2c(\lambda)\epsilon^{\prime}(\lambda).$
Then $h_{3}\leq h_{2}\leq h_{1}$, $\liminf_{\lambda\to\infty}h_{i}(\lambda)>0$ and $\limsup_{\lambda\to 1+}h_{i}(\lambda)<0$ for all $i$.
Proposition 8 follows from properties listed in Theorem 21, Lemma 29, 30 and 31. If we consider an affine cost function $c(\lambda)=a+b\lambda$, then Proposition 8 implies that if we used $G=h_{3}$ as the gradient like function, the gradient descent is expected to remain bounded, and therefore the stochastic gradient descent is expected to converge to a stationary point within $(1,\infty)$. Furthermore, the minimiser of $\tilde{L}(\lambda)$ corresponds to the zero of $h_{2}$, which has similar behaviour. We conclude this section by a (crude) upper bound for the minimiser of true asymptotic variance:
Remark 9.
For a cost function $c(\lambda)=a+b\lambda$ where $a\geq 0$ and $b>0$ and a test function $f\in L^{2}(\pi)$, $\mathrm{var}_{\pi}(f)>0$, the minimiser $\lambda^{*}$ of $c(\lambda)\mathrm{var}(P_{\lambda},f)$ satisfies $2\leq\lambda^{*}\leq 4\sqrt{\hat{w}(a/b+1)}+4\hat{w}+1$ by Proposition 51 and 54 in Appendix F.
6. Convexity and monotonicity of i-SIR covariances and acceptance rates
Next, we discuss some technical properties of the i-SIR (Algorithm 1) and its continuous relaxation (Algorithm 2). Namely, Sections 6.1 and 6.2 establish properties for asymptotic covariances and rejection probability for the i-SIR, and Section 6.3 extends these properties to the continuous relaxation. In particular, the results are necessary for establishing convexity of asymptotic variance in Section 7. Recall the notion of convexity of a sequence, or sequential convexity [cf. mitrinovic].
Definition 10.
A real sequence $(a_{i})$ is called convex if $2a_{i+1}\leq a_{i}+a_{i+2}$ for all $i$.
Sequential convexity is convenient because it ensures convexity for the interpolations of the sequence; see Lemma 44 (i) in Appendix D.
We will now give the definitions we require. We denote the natural numbers as $\mathbb{N}=\{0,1,2,\dots\}$ and positive natural numbers as $\mathbb{N}_{+}=\{1,2,\dots\}$. Throughout the upcoming sections we have fixed $\pi$ and $q$ as mentioned in Section 1 (more detailed in Appendix A). We use the shorthand notation $\pi(\mathrm{d}x)=\pi(x)\mathrm{d}\mu(x)$ and $q(\mathrm{d}x)=q(x)\mathrm{d}\mu(x)$, where $\mu$ is the dominating measure. When we speak of atoms, we mean atoms with respect to $\pi$. The support of $\pi$ is denoted by $\mathbb{S}$, the set of atoms in the support by $\mathbb{S}_{+}=\{x\in\mathbb{S}:\pi(\{x\})>0\}$ and the set of non-atoms by $\mathbb{S}_{0}=\{x\in\mathbb{S}:\pi(\{x\})=0\}$. We define the inner product $\langle g|h\rangle_{\pi}=\int g(x)h(x)\pi(\mathrm{d}x)$ for $g,h\in L^{2}(\pi)$, the corresponding norm $\|f\|_{\pi}=\sqrt{\langle f|f\rangle_{\pi}}$, and the function spaces
$L_{0}^{2}(\pi)=\{g\in L^{2}(\pi):\mathbb{E}_{\pi}[g]=0\},\qquad L_{0,+}^{2}(\pi)=\{g\in L_{0}^{2}(\pi):\|g\|_{\pi}>0\}.$
The centred version of a $\pi$-integrable function is denoted as $\bar{f}=f-\mathbb{E}_{\pi}[f]$. The counting measure, that is, the number of elements in a set, is denoted by #\#.
In our proofs, (permutation) symmetricity of functions plays a central role, because of exchangeability of the variables.
Definition 11.
Let $\in\mathbb{N}_{+}, m<nm<n$, $U\subset\mathbb{X}^{n}$ and $f:U\to\mathbb{R}$. The function is called symmetric if
$f(x_{1},\dots,x_{n})=f(x_{i_{1}},\dots,x_{i_{n}})$
for any permutation $i_{1},\dots,i_{n}$ of $1,\dots,n$, and symmetric with respect to the first $m$ arguments if
$f(x_{1},\dots,x_{m},x_{m+1},\dots x_{n})=f(x_{i_{1}},\dots,x_{i_{m}},x_{m+1},\dots x_{n})$
for any permutation $i_{1},\dots,i_{m}$ of $1,\dots,m$.
Note that in Definition 11 the arguments $x_{i}\in\mathbb{X}$. Let $N\in\mathbb{N}_{+}$, then in particular the mapping $z\mapsto\sum_{i=1}^{N}w(z_{i})$ is symmetric. We use the following notation for set of states where this mapping is positive:
$\mathbb{W}_{N}=\bigg\{z\in\mathbb{X}^{N}\;:\;\sum_{i=1}^{N}w(z_{i})>0\bigg\}=\bigcup_{i=1}^{N}\bigg\{z\in\mathbb{X}^{N}\;:\;w(z_{i})>0\bigg\}.$
We will also use the following notation
$\int_{\mathbb{X}^{N-1}}f(x)\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)=f(x),$
when $N=1$. When we use a test set $A$ we always mean a measurable set. The transition probability which we obtain from Algorithm 1 when $N\in\mathbb{N}_{+}$ will be denoted by $P_{N}$ throughout the upcoming sections, and is defined in (1),
$P_{N}(z_{1},\mathrm{d}y)=\int_{\mathbb{X}^{N-1}}\sum_{i=1}^{N}\frac{w(z_{i})}{\sum_{j=1}^{N}w(z_{j})}\delta_{z_{i}}(\mathrm{d}y)\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
with $P_{1}(z_{1},\mathrm{d}y)=\delta_{z_{1}}(\mathrm{d}y)$.
6.1. Stationary covariances
If $(X_{k})_{k\geq 1}$ is the stationary i-SIR chain, then the expressions $\mathbb{E}[f(X_{k})f(X_{k+1})]$ may be written in terms of inner products $\langle f|P_{N}f\rangle_{\pi}$. In particular, when $\pi(f)=0$, this expression coincides with $\mathrm{Cov}\big(f(X_{k}),f(X_{k+1})\big)$, so we will call these inner products ‘covariances’ hereafter.
For $f\in L^{2}(\pi)$ and $N\in\mathbb{N}_{+}$, we have $P_{N}f\in L^{2}(\pi)$ and therefore the inner products are well-defined and finite. They can be written as
(6) $\begin{split}\langle f|P_{N}f\rangle_{\pi}&=\int_{\mathbb{S}}\int_{\mathbb{X}^{N-1}}\sum_{i=1}^{N}\frac{w(z_{i})f(z_{1})f(z_{i})}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\pi(\mathrm{d}z_{1})\\ &=\int_{\mathbb{W}_{N}}\sum_{i=1}^{N}\frac{w(z_{1})w(z_{i})f(z_{1})f(z_{i})}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N}q(\mathrm{d}z_{n})\bigg),\end{split}$
because $w(z_{1})q(\mathrm{d}z_{1})=\pi(\mathrm{d}z_{1})$. Note also that $\langle f|P_{1}f\rangle=\langle f|f\rangle_{\pi}$. The following result shows that integrals similar to the one above can be written in a form involving a quadratic term:
Lemma 12.
Let $N,M\in\mathbb{N}_{+}$, $f\in L^{2}(\pi)$ and $h:\mathbb{W}_{N}\times M\to[0,\infty)$ be measurable bounded function and symmetric with respect to the first $N$ arguments, then
$\int_{\mathbb{W}_{N}\times\mathbb{X}^{M}}h(z)\sum_{i=1}^{N}\frac{w(z_{1})w(z_{i})f(z_{1})f(z_{i})}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N+M}q(\mathrm{d}z_{n})\bigg)$
$\quad=\int_{\mathbb{W}_{N}\times\mathbb{X}^{M}}h(z)\frac{1}{N}\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N+M}q(\mathrm{d}z_{n})\bigg).$
Proof.
Since $h$ is symmetric with respect to the first $N$ arguments and $z_{1},\dots,z_{N}$ are exchangeable under the integral, the expression on the left can be written as
$\int_{\mathbb{W}_{N}\times\mathbb{X}^{M}}h(z)\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{N}\frac{w(z_{i})w(z_{k})f(z_{i})f(z_{k})}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N+M}q(\mathrm{d}z_{n})\bigg),$
which equals the expression on the right. ∎
Lemma 12 shows that the covariances in equation (6) are positive:
Lemma 13.
Let $N\in\mathbb{N}_{+}$ and $f\in L^{2}(\pi)$, then
$\langle f|P_{N}f\rangle_{\pi}=\frac{1}{N}\int_{\mathbb{W}_{N}}\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N}q(\mathrm{d}z_{n})\bigg)\geq 0,$
and $\langle f|P_{N}f\rangle_{\pi}=0$ if and only if $\|f\|_{\pi}=0$.
Proof.
Equation follows from Lemma 12 for any $M\in\mathbb{N}_{+}$ and $h=1$. If $\|f\|_{\pi}=0$, then $f=0$ $\pi$-almost surely, hence (6) implies $\langle f|P_{N}f\rangle_{\pi}=0$. If $\|f\|_{\pi}\neq 0$, then $f\neq 0$ in a $\pi$-positive set and thus at least $A=\{x\in\mathbb{S}:f(x)>0\}$ or $B=\{x\in\mathbb{S}:f(x)<0\}$ has to be $\pi$-positive, which is also $q$-positive. It holds for all $x\in\mathbb{S}$ that $w(x)>0$, hence if $A$ is $\pi$-positive it follows that
$\langle f|P_{N}f\rangle_{\pi}\geq\frac{1}{N}\int_{A^{N}}\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N}q(\mathrm{d}z_{n})\bigg)>0$
and if $B$ is $\pi$-positive, we get similar inequality. ∎
The following lemma connects expressions similar to the one above for different $N$:
Lemma 14.
Let $N\in\mathbb{N}_{+}$, $f\in L^{2}(\pi)$ and let $h:\mathbb{W}_{N+1}\times\mathbb{X}\to[0,\infty)$ be measurable bounded function and symmetric with respect to the first $N+1$ arguments, then
(7) $\begin{split}&\frac{1}{N}\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}h(z)\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)\\ &\quad\geq\frac{1}{N+1}\int_{\mathbb{W}_{N+1}\times\mathbb{X}}h(z)\frac{\Big(\sum_{i=1}^{N+1}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{j=1}^{N+1}w(z_{j})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).\end{split}$
If $\bar{f}\in L_{0,+}^{2}(\pi)$ and there exists a $\pi$-positive set $A$ such that $h(z)>0$ for every $z\in\mathbb{W}_{N+1}\times A$, then the inequality is strict.
Proof.
Let us define $g_{k}:\mathbb{W}_{N+1}\times\mathbb{X}\to\mathbb{R}$,
$g_{k}(z_{1},\dots,z_{N+2})=\begin{cases}\frac{\big(\sum_{i=1,i\neq k}^{N+1}w(z_{i})f(z_{i})\big)^{2}}{\sum_{j=1,j\neq k}^{N+1}w(z_{j})},&\sum_{l=1,l\neq k}^{N+1}w(z_{l})>0\\ 0,&\sum_{l=1,l\neq k}^{N+1}w(z_{l})=0,\end{cases}$
for $k\in\{1,2,\dots,N+2\}$. We have
(8) $\int_{B}\frac{h(z)}{(N+1)N}\sum_{k=1}^{N+1}g_{k}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)\geq\int_{B}\frac{h(z)}{N+1}g_{N+2}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
for any $B\subset\mathbb{W}_{N+1}\times\mathbb{X}$ that is measurable with respect to the product sigma-algebra by Lemma 41 in Appendix C. When $B=\mathbb{W}_{N+1}\times\mathbb{X}$, the right-hand side of (8) equals the right-hand side of (7), and
$\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{h(z)}{(N+1)N}\sum_{k=1}^{N+1}g_{k}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)=\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{h(z)}{N}g_{N+1}(z)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\quad=\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}\frac{h(z)}{N}\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{j=1}^{N}w(z_{j})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg),$
where the first equality follows since $h$ is symmetric with respect to the first $N+1$ arguments and $z_{1},\dots,z_{N+1}$ are exchangeable under the integral, and the second from $\mathbb{W}_{N}\times\mathbb{X}\subset\mathbb{W}_{N+1}$.
If $\bar{f}\in L_{0,+}^{2}(\pi)$, then there exist two disjoint $\pi$-positive sets $U,V\subset\mathbb{S}$ such that $f$ differs on them, these sets and set AA in the assumption are $q$-positive. Let $W=U\times V\times\mathbb{S}^{N-1}\times A$ (for $N=1$, $W=U\times V\times A$). If $B=W$, we get strict inequality for (8) by Lemma 41 and 42 in Appendix C, thus strict inequality also holds when $B=\mathbb{W}_{N+1}\times\mathbb{X}=W\cup((\mathbb{W}_{N+1}\times\mathbb{X})\setminus W)$, from which the claim follows. ∎
In particular, Lemma 14 implies that covariance decreases with respect to $N$:
Lemma 15.
Let $N\in\mathbb{N}_{+}$ and $f\in L^{2}(\pi)$, then
$\langle f|P_{N}f\rangle_{\pi}\geq\langle f|P_{N+1}f\rangle_{\pi},$
where the inequality is strict if and only if $\bar{f}\in L_{0,+}^{2}(\pi)$.
Proof.
If $\bar{f}\notin L_{0,+}^{2}(\pi)$, then there exists $c\in\mathbb{R}$ such that $f=c$ $\pi$-almost surely, hence $\langle f|P_{N}f\rangle_{\pi}=c^{2}$ for all $N\in\mathbb{N}_{+}$. Combining this with Lemma 14 with $h=1$ we obtain the claim. ∎
The next result shows that the covariances form a convex sequence with respect to $N$:
Lemma 16.
Let $N\in\mathbb{N}_{+}$ and $f\in L^{2}(\pi)$, then
$\langle f|P_{N}f\rangle-2\langle f|P_{N+1}f\rangle+\langle f|P_{N+2}f\rangle\geq 0,$
where the inequality is strict if and only if $\bar{f}\in L_{0,+}^{2}(\pi)$.
Proof.
We first show that
$\langle f|P_{N}f\rangle_{\pi}-2\langle f|P_{N+1}f\rangle_{\pi}+\langle f|P_{N+2}f\rangle_{\pi}$
(9) $\quad=2\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\sum_{i=1}^{N}w(z_{1})f(z_{1})w(z_{i})f(z_{i})}{\sum_{k=1}^{N}w(z_{k})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\qquad-2\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\sum_{i=1}^{N+1}w(z_{1})f(z_{1})w(z_{i})f(z_{i})}{\sum_{k=1}^{N+1}w(z_{k})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
Let us denote $A_{n}=\sum_{i=1}^{n}w(z_{1})w(z_{i})f(z_{1})f(z_{i})$ and $B_{n}=\sum_{i=1}^{n}w(z_{i})$. We have
$\langle f|P_{N}f\rangle-2\langle f|P_{N+1}f\rangle-\langle f|P_{N+2}f\rangle$
$\quad=\int_{\mathbb{S}\times\mathbb{X}^{N+1}}\Big(\frac{A_{N}}{B_{N}}-2\frac{A_{N+1}}{B_{N+1}}+\frac{A_{N+2}}{B_{N+2}}\Big)\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\quad=\int_{\mathbb{S}\times\mathbb{X}^{N+1}}\frac{B_{N+2}B_{N+1}A_{N}-2B_{N+2}B_{N}A_{N+1}+B_{N+1}B_{N}A_{N+2}}{B_{N+2}B_{N+1}B_{N}}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
By Lemma 43 in Appendix C.
$B_{N+2}B_{N+1}A_{N}-2B_{N+2}B_{N}A_{N+1}+B_{N+1}B_{N}A_{N+2}$
$\quad=2w(z_{N+2})B_{N+1}A_{N}-2w(z_{N+2})B_{N}A_{N+1}$
$\qquad-B_{N+1}B_{N}w(z_{1})f(z_{1})[w(z_{N+1})f(z_{N+1})-w(z_{N+2})f(z_{N+2})]$
$\qquad+B_{N+1}A_{N}[w(z_{N+1})-w(z_{N+2})].$
The claim follows since $z_{N+1}$ and $z_{N+2}$ are exchangeable under the following integrals
$\int_{\mathbb{S}\times\mathbb{X}^{N+1}}\frac{w(z_{1})f(z_{1})w(z_{N+1})f(z_{N+1})}{B_{N+2}}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\quad=\int_{\mathbb{S}\times\mathbb{X}^{N+1}}\frac{w(z_{1})f(z_{1})w(z_{N+2})f(z_{N+2})}{B_{N+2}}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
and
$\int_{\mathbb{S}\times\mathbb{X}^{N+1}}\frac{A_{N}w(z_{N+1})}{B_{N+2}B_{N}}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)=\int_{\mathbb{S}\times\mathbb{X}^{N+1}}\frac{A_{N}w(z_{N+2})}{B_{N+2}B_{N}}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
Combining (9) with Lemma 12, it follows that $\langle f|P_{N}f\rangle_{\pi}-2\langle f|P_{N+1}f\rangle_{\pi}+\langle f|P_{N+2}f\rangle_{\pi}$ may be written as
$\frac{2}{N}\int_{\mathbb{W}_{N}\times\mathbb{X}^{2}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\Big(\sum_{i=1}^{N}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{k=1}^{N}w(z_{k})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\qquad-\frac{2}{N+1}\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{w(z_{N+2})}{\sum_{j=1}^{N+2}w(z_{j})}\frac{\Big(\sum_{i=1}^{N+1}w(z_{i})f(z_{i})\Big)^{2}}{\sum_{k=1}^{N+1}w(z_{k})}\bigg(\prod_{n=1}^{N+2}q(\mathrm{d}z_{n})\bigg).$
Therefore, the claim follows from Lemma 14 and 15. ∎
Covariances can be expressed with the help of i.i.d. random variables from $q$. We will use this property and the fact that if $X\sim q$, then $\mathbb{E}[w(X)f(X)]=\mathbb{E}_{\pi}[f]$ for $\pi$-integrable $f$ and $\mathbb{E}[w(X)]=1$, to show the following limit property:
Proposition 17.
Let $f\in L^{2}(\pi)$, then
$\langle f|P_{N}f\rangle\xrightarrow{N\to\infty}\mathbb{E}_{\pi}[f]^{2}.$
Proof.
Let $Z_{1},Z_{2},\dots$ be i.i.d. from $q$. Let us define sequences of random variables $(A_{N})$, $(B_{N})$, $(C_{N})$ and $(D_{N})$:
$A_{N}$ $=\frac{1}{N}\frac{\sum_{i=1}^{N}w(Z_{i})^{2}f(Z_{i})^{2}}{\sum_{j=1}^{N}w(Z_{j})},\qquad$ $B_{N}$ $=\frac{1}{N}\frac{\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}w(Z_{i})w(Z_{j})f(Z_{i})f(Z_{j})}{\sum_{k=1}^{N}w(Z_{k})},$
$C_{N}$ $=\frac{w(Z_{1})^{2}f(Z_{1})^{2}}{\sum_{j=1}^{N}w(Z_{j})},\qquad$ $D_{N}$ $=\frac{1}{N^{2}}\frac{\big(\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}w(Z_{i})w(Z_{j})f(Z_{i})f(Z_{j})\big)^{2}}{\sum_{k=1}^{N}\sum_{l=1,l\neq k}^{N}w(Z_{k})w(Z_{l})},$
whenever the denominator is positive, and otherwise 0. We have by Lemma 13, that
$\langle f|P_{N}f\rangle=\mathbb{E}[A_{N}]+\mathbb{E}[B_{N}].$
By strong law of large numbers $C_{N}\to 0$ and $A_{N}+B_{N}\to\mathbb{E}_{\pi}[f]^{2}$ almost surely. Since $|C_{N}(Z)|\leq w(Z_{1})f(Z_{1})^{2}$ for all $N$, we have by dominated convergence $\mathbb{E}[C_{N}]\to 0$ and because $Z_{i}$ are exchangeable, we have $\mathbb{E}[A_{N}]=\mathbb{E}[C_{N}]\to 0$, hence $A_{N}\to 0$ in probability and we obtain $B_{N}\to\mathbb{E}_{\pi}[f]^{2}$ in probability. It holds that $\mathbb{E}[|B_{N}|^{2}]\leq\mathbb{E}[D_{N}]$ and for $N\geq 2$ we have
$\mathbb{E}[D_{N}]\leq\mathbb{E}\Big[\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}w(Z_{i})w(Z_{j})f(Z_{i})^{2}f(Z_{j})^{2}\Big]=\frac{N(N-1)}{N^{2}}\mathbb{E}_{\pi}[f^{2}]^{2},$
where the inequality follows from Sedrakyan’s inequality (Lemma 40 in Appendix C) and the equality from the fact that $Z_{1},\dots,Z_{N}$ are i.i.d. Thus, $\sup_{N}\mathbb{E}[|B_{N}|^{2}]<\infty$, which implies that $(B_{N})$ is uniformly integrable and therefore $\mathbb{E}[B_{N}]\to E_{\pi}[f]^{2}$. ∎
6.2. Rejection probabilities
The rejection probability for Algorithm 1 is defined as
$\epsilon(N,z_{1})=\int_{X^{N-1}}\frac{w(z_{1})}{\sum_{i=1}^{N}w(z_{i})}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
and the average rejection probability
$\epsilon(N)=\int_{\mathbb{S}}\epsilon(N,z_{1})\pi(\mathrm{d}z_{1})=\int_{\mathbb{W}_{N}}\frac{w(z_{1})^{2}}{\sum_{i=1}^{N}w(z_{i})}\bigg(\prod_{n=1}^{N}q(\mathrm{d}z_{n})\bigg),$
when $N\in\mathbb{N}_{+}$. We shall use these notations throughout the upcoming sections. Monotonicity and sequential convexity are quite easy to show for the rejection probabilities:
Lemma 18.
Let $N\in\mathbb{N}_{+}$ and $x\in\mathbb{S}$, then
$\epsilon(N)-\epsilon(N+1)$ $>0,\qquad$ $\epsilon(N)-2\epsilon(N+1)+\epsilon(N+2)$ $>0,$
$\epsilon(N,x)-\epsilon(N+1,x)$ $>0,\qquad$ $\epsilon(N,x)-2\epsilon(N+1,x)+\epsilon(N+2,x)$ $>0.$
Proof.
Let us denote $z_{1}=x$ and $B_{n}=\sum_{i=1}^{n}w(z_{i})$, then
$\epsilon(N,z_{1})-\epsilon(N+1,z_{1})$ $=w(z_{1})\int_{\mathbb{X}^{N}}\frac{w(z_{N+1})}{B_{N}B_{N+1}}\bigg(\prod_{n=2}^{N+1}q(\mathrm{d}z_{n})\bigg)>0.$
We have also
$\epsilon(N,z_{1})-2\epsilon(N+1,z_{1})+\epsilon(N+2,z_{1})$
$\quad=w(z_{1})\int_{\mathbb{X}^{N+1}}\frac{B_{N+1}B_{N+2}-2B_{N}B_{N+2}+B_{N}B_{N+1}}{B_{N}B_{N+1}B_{N+2}}\bigg(\prod_{n=2}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\quad=w(z_{1})\int_{\mathbb{X}^{N+1}}\frac{[w(z_{N+1})-w(z_{N+2})]B_{N+1}+2w(z_{N+1})w(z_{N+2})}{B_{N}B_{N+1}B_{N+2}}\bigg(\prod_{n=2}^{N+2}q(\mathrm{d}z_{n})\bigg)$
$\quad=w(z_{1})\int_{\mathbb{X}^{N+1}}\frac{2w(z_{N+1})w(z_{N+2})}{B_{N}B_{N+1}B_{N+2}}\bigg(\prod_{n=2}^{N+2}q(\mathrm{d}z_{n})\bigg)>0,$
where the second equality follows from Lemma 43 in Appendix C, and the third from
$w(z_{1})\int_{\mathbb{X}^{N+1}}\frac{w(z_{N+1})}{B_{N}B_{N+2}}\bigg(\prod_{n=2}^{N+2}q(\mathrm{d}z_{n})\bigg)=w(z_{1})\int_{\mathbb{X}^{N+1}}\frac{w(z_{N+2})}{B_{N}B_{N+2}}\bigg(\prod_{n=2}^{N+2}q(\mathrm{d}z_{n})\bigg),$
where the equality follows since $z_{N+1}$ and $z_{N+2}$ are exchangeable under the integral. Strict inequalities hold also for the average rejection probability, since $\epsilon(N)=\int_{\mathbb{S}}\epsilon(N,x)\pi(\mathrm{d}x)$. ∎
Note that the holding probability of Algorithm 1 for $z_{1}\in\mathbb{S}$ satisfies
$P_{N}(z_{1},\{z_{1}\})$ $=\int_{\mathbb{X}^{N-1}}\sum_{i=1}^{N}\frac{w(z_{i})}{\sum_{k=1}^{N}w(z_{k})}\mathbf{1}\left(z_{i}=z_{1}\right)\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg).$
In the case $z_{1}$ is not an atom, the holding probability $P_{N}(z_{1},\{z_{1}\})$ coincides with the rejection probability $\epsilon(N,z_{1})$, for which the monotonicity and sequential convexity are shown above. When $z_{1}$ is an atom, we cannot use the same results directly, but we can show instead that there exists $g\in L^{2}(\pi)$ such that $P_{N}(z_{1},\{z_{1}\})=\langle g|P_{\lambda}g\rangle_{\pi}$ (see Lemma 22 in Section 6.3), so the same properties follow from results in Section 6.1.
Rejection probability and holding probability can be expressed with the help of i.i.d. random variables from $q$. We will use this property to show the following lemma:
Lemma 19.
Let $x\in\mathbb{S}$. It holds that
$\lim_{N\to\infty}\epsilon(N)=0,\qquad\lim_{N\to\infty}\epsilon(N,x)=0,\qquad\lim_{N\to\infty}P_{N}(x,\{x\})=\pi(\{x\}).$
Proof.
Let $Z_{2},Z_{3},\dots$ be i.i.d. from $q$, then
$P_{N}(x,\{x\})=\mathbb{E}\bigg[\frac{w(x)+\sum_{i=2}^{N}w(Z_{i})\mathbf{1}\left(Z_{i}=x\right)}{w(x)+\sum_{j=2}^{N}w(Z_{j})}\bigg],\qquad\epsilon(N,x)=\mathbb{E}\bigg[\frac{w(x)}{w(x)+\sum_{j=2}^{N}w(Z_{j})}\bigg].$
In both equations the term inside the expectation is dominated by $1$ for all $N$ and by strong law of large numbers the term in the first equation converges to $\pi(\{x\})$ and term in the second equation converges to 0 almost surely as $N\to\infty$. Therefore, by dominated convergence we obtain the claim for holding probability and rejection probability. Thus, the claim for average rejection probability follows by dominated convergence. ∎
6.3. Properties for the continuous relaxation
We can generalise the properties of the i-SIR (Algorithm 1) for its continuous relaxation (Algorithm 2). This is possible, since as mentioned in Section 2 it holds for the Markov chain of Algorithm 2, that if $\lambda\in[1,\infty)$, then
$P_{\lambda}(x,A)=\beta(\lambda)P_{\lfloor\lambda\rfloor}(x,A)+\big(1-\beta(\lambda)\big)P_{\lfloor\lambda\rfloor+1}(x,A),$
where $\beta(\lambda)=\lfloor\lambda\rfloor+1-\lambda\in(0,1]$, which yields
(10) $\begin{split}\langle f|P_{\lambda}f\rangle_{\pi}&=\beta(\lambda)\langle f|P_{\lfloor\lambda\rfloor}f\rangle_{\pi}+\big(1-\beta(\lambda)\big)\langle f|P_{\lfloor\lambda\rfloor+1}f\rangle_{\pi}\\ \epsilon(\lambda,x)&=\beta(\lambda)\epsilon(\lfloor\lambda\rfloor,x)+\big(1-\beta(\lambda)\big)\epsilon(\lfloor\lambda\rfloor+1,x)\\ \epsilon(\lambda)&=\beta(\lambda)\epsilon(\lfloor\lambda\rfloor)+\big(1-\beta(\lambda)\big)\epsilon(\lfloor\lambda\rfloor+1),\end{split}$
where $f\in L^{2}(\pi)$ and $x\in\mathbb{S}$. We use this notation throughout the upcoming sections. Because of the interpolation form of (10), the properties established for integer $\lambda$ above can be generalised to the space $1,\infty)$:
Theorem 20.
Let $f\in L^{2}(\pi)$, then $\lambda\mapsto\langle f|P_{\lambda}f\rangle_{\pi}$, $\lambda\geq 1$ is continuous non-negative decreasing convex function. Additionally, we have the following properties:
  1. (i)
    For any $\lambda\geq 1$ it holds that $\langle f|P_{\lambda}f\rangle_{\pi}=0$ if and only if $\|f\|_{\pi}=0$.
  2. (ii)
    $\lambda\mapsto\langle f|P_{\lambda}f\rangle_{\pi}$ is strictly decreasing if and only if $\bar{f}\in L_{0,+}^{2}(\pi)$.
Proof.
Follows from Lemma 44 in Appendix D when combined with Lemma 13, 15 and 16. ∎
Theorem 21.
It holds that $\lambda\mapsto\epsilon(\lambda)$, $\lambda\geq 1$ is positive continuous convex and strictly decreasing function, same holds for $\lambda\mapsto\epsilon(\lambda,x)$, $\lambda\geq 1$ when $x\in\mathbb{S}$.
Proof.
By definition $\epsilon(N)$ and $\epsilon(N,x)$ are positive for all $N\in\mathbb{N}_{+}$. Therefore, the claim follows from Lemma 44 in Appendix D when combined with Lemma 18. ∎
Lemma 22.
Let $\lambda\geq 1$ and $x\in\mathbb{S}$. If $x$ is an atom, then
$P_{\lambda}(x,\{x\})=\langle g_{x}|P_{\lambda}g_{x}\rangle_{\pi},$
where $g_{x}(y)=\frac{1}{\sqrt{\pi(\{x\})}}\mathbf{1}\left(y=x\right)$.
Proof.
Since $x$ is an atom, we have $P_{1}(x,\{x\})=\langle g_{x}|g_{x}\rangle_{\pi}$ and if $N\in\mathbb{N}_{+}$, $N\geq 2$, then
$P_{N}(x,\{x\})$ $=\int_{\mathbb{X}^{N-1}}\frac{w(x)+\sum_{i=2}^{N}w(z_{i})\mathbf{1}\left(z_{i}=x\right)}{w(x)+\sum_{k=2}^{N}w(z_{k})}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
$=\int_{\mathbb{S}}\int_{\mathbb{X}^{N-1}}\sum_{i=1}^{N}\frac{w(z_{i})\mathbf{1}\left(z_{i}=x\right)}{\sum_{k=1}^{N}w(z_{k})}\frac{\mathbf{1}\left(z_{1}=x\right)}{\pi(\{x\})}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\pi(\mathrm{d}z_{1}).$
Thus, by (6) we get the claim for $N\in\mathbb{N}_{+}$, from which the general version follows. ∎
Theorem 23.
If $(\mathbb{S})>1$ and $x\in\mathbb{S}$, then $\lambda\mapsto P_{\lambda}(x,\{x\})$, $\lambda\geq 1$ is positive continuous convex and strictly decreasing function.
Proof.
When $x$ is not an atom, the claim follows from Theorem 21. When $x$ is an atom, the claim follows from Lemma 22 and Theorem 20 by noticing $\bar{g}_{x}\in L_{0,+}^{2}(\pi)$. ∎
We will also give convexity results for the functions related to the approximate i-SIR.
Lemma 24.
The function
$\lambda\mapsto\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)},\qquad\lambda>1$
is positive, strictly convex and decreasing.
Proof.
Function $\epsilon:(1,\infty)\to[0,1)$, $\epsilon(\lambda)$ is positive convex and strictly decreasing by Theorem 21. Thus, by noticing that $g:[0,1)\to\mathbb{R}$, $g(x)=\frac{1+x}{1-x}$ is positive strictly increasing strictly convex function, we get that $g(\epsilon(\lambda))$ is positive strictly decreasing function and it holds for $1<\lambda_{1}<\lambda_{2}$ and $t\in(0,1)$ that
$g\big(\epsilon(t\lambda_{1}+(1-t)\lambda_{2})\big)\leq g\big(t\epsilon(\lambda_{1})+(1-t)\epsilon(\lambda_{2})\big)<tg\big(\epsilon(\lambda_{1})\big)+(1-t)g\big(\epsilon(\lambda_{2})\big).\qed$
When integrating holding probability, we can divide the integral into two parts
$\int_{\mathbb{S}}P_{\lambda}(y,\{y\})\pi(\mathrm{d}y)$ $=\int_{\mathbb{S}_{0}}\epsilon(\lambda,y)\pi(\mathrm{d}y)+\int_{\mathbb{S}_{+}}P_{\lambda}(y,\{y\})\pi(\mathrm{d}y),$
$\int_{\mathbb{S}}\pi(\{x\})\pi(\mathrm{d}x)$ $=\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x).$
The equations above are helpful in the following lemma:
Lemma 25.
If $(\mathbb{S})>1$, then
$\lambda\mapsto\frac{\int_{\mathbb{S}}[P_{\lambda}(x,\{x\})-\pi(\{x\})]\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}}\pi(\{x\})\pi(x)},\qquad\lambda\geq 1$
is positive convex and strictly decreasing function and
$\lambda\mapsto\frac{1+\psi(\lambda)}{1-\psi(\lambda)},\qquad\lambda>1$
is positive strictly convex and decreasing function.
Proof.
Function $\lambda\mapsto\int_{\mathbb{S}_{0}}\epsilon(\lambda,x)\pi(\mathrm{d}x)+\int_{\mathbb{S}_{+}}\pi(\{x\})P_{\lambda}(x,\{x\})\in(\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x),1]$ is positive convex and strictly decreasing by Theorem 21, 23 and Lemma 19. Therefore, we obtain the claim for $\psi$ by noticing $g:(\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x),1]\to(0,1]$, $g(y)=\frac{y-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}$ is linear function. Since $\psi(1)=1$, we have $\psi(y)\in[0,1)$ for all $y\in(1,\infty)$, hence rest of the proof is similar to the proof of Lemma 24. ∎
7. Asymptotic variance
One of our main results is showing for Algorithm 2, that the asymptotic variance as a function $\lambda\mapsto\mathrm{var}(P_{\lambda},f)$, $\lambda\in[1,\infty)$ is monotonic and convex for all $f\in L^{2}(\pi)$. The monotonicity can sometimes be shown with the Peskun order [peskun, tierney-note], but for i-SIR, a straightforward calculation shows a contradiction with
$\pi(x)=\begin{cases}2x,&x\in[0,1]\\ 0,&x\notin[0,1],\end{cases}\qquad q(x)=\begin{cases}1,&x\in[0,1]\\ 0,&x\notin[0,1],\end{cases}$
when choosing $x=0.2$ and comparing $P_{2}(x,\;\cdot\;)$ and $P_{3}(x,\;\cdot\;)$ with both sets $(0,0.1)$ and $(0.9,1)$. However, there is a weaker condition for showing monotonicity of asymptotic variance, namely the covariance ordering [tierney-note, mira-geyer], which in our setting means the monotonicity of $\lambda\mapsto\langle f|P_{\lambda}f\rangle_{\pi}$ for all $f\in L^{2}(\pi)$. In this paper monotonicity and convexity of asymptotic variance are obtained in a similar fashion as in [e.g. andrieu-vihola-order, Proposition 26] using Dirichlet forms, these results are combined in Theorem 63 in Appendix H. Without loss of generality we can study $L_{0}^{2}(\pi)$ functions.
Lemma 26.
Let $\lambda\geq 1$, then $P_{\lambda}$ is $\pi$-reversible. If Assumption (A1) holds, then $I-P_{\lambda}:L_{0}^{2}(\pi)\to L_{0}^{2}(\pi)$ is invertible for $\lambda>1$.
Proof.
We have $P_{N}$ is $\pi$-reversible for $N\in\mathbb{N}_{+}$, $N\geq 2$ [e.g. tjelmeland, andrieu-lee-vihola], which also holds for $N=1$, since $P_{1}(x,\mathrm{d}y)=\delta_{x}(\mathrm{d}y)$. Therefore, we have $P_{\lambda}$ is $\pi$-reversible for $\lambda\geq 1$, since
$P_{\lambda}(x,\mathrm{d}y)=(\lfloor\lambda\rfloor+1-\lambda)P_{\lfloor\lambda\rfloor}(x,\mathrm{d}y)+(\lambda-\lfloor\lambda\rfloor)P_{\lfloor\lambda\rfloor+1}(x,\mathrm{d}y),$
If $\lambda>1$, then by Lemma 49 in Appendix E
$P_{\lambda}(x,A)=c\pi(A)+(1-c)Q(x,A),$
where $c=\frac{\lambda-1}{2\hat{w}+\lambda-1}$ and $Q$ is $\pi$-reversible Markov kernel. Thus, for $f\in L_{0}^{2}(\pi)$
$(I-P_{\lambda})f=f-(1-c)Qf=cf+(1-c)(I-Q)f$
and
$\langle(I-P_{\lambda})f|(I-P_{\lambda})f\rangle_{\pi}$ $=c^{2}\langle f|f\rangle_{\pi}+c(1-c)\langle f|(I-Q)f\rangle_{\pi}+c(1-c)\langle(I-Q)f|f\rangle_{\pi}$
$\quad+(1-c)^{2}\langle(I-Q)f|(I-Q)f\rangle_{\pi}$
$\geq c^{2}\langle f|f\rangle_{\pi},$
since the Dirichlet form is non-negative for Markov kernel. Therefore, $I-P_{\lambda}$ is invertible by Lemma 59 in Appendix H. ∎
Theorem 27.
Let Assumption (A1) hold and $f\in L_{0,+}^{2}(\pi)$, then $\lambda\mapsto\mathrm{var}(P_{\lambda},f)$, $\lambda>1$ is strictly convex and decreasing function.
Proof.
The claim follows from Theorem 63 (v) in Appendix H when combined with Lemma 26 and Theorem 20. ∎
In some applications Dirichlet form can be more interesting than the inner product $\langle f|P_{\lambda}f\rangle_{\pi}$. Let $\lambda\geq 1$, then the Dirichlet form for Algorithm 2 is defined
$\mathcal{E}_{\lambda}(f)=\langle f|(I-P_{\lambda}f)\rangle_{\pi},\qquad$
where $f\in L^{2}(\pi)$. Thus, we can give another version of Theorem 20:
Proposition 28.
Let $f\in L_{0,+}^{2}(\pi)$, then for Algorithm 2 it holds that $\lambda\mapsto\mathcal{E}_{\lambda}(f)$, $\lambda\geq 1$ is non-negative, continuous, concave and strictly increasing function.
Proof.
Since $\mathcal{E}_{\lambda}(g)=\langle f|f\rangle_{\pi}-\langle f|P_{\lambda}f\rangle_{\pi}$ and $\langle f|P_{1}f\rangle_{\pi}=\langle f|f\rangle_{\pi}$, the claim follows from Theorem 20. ∎
8. Bounds for i-SIR rejection rate and asymptotic variance
We shall now present upper and lower bounds for certain functions obtained from Algorithm 2. These bounds may also provide asymptotic speed of the function.
8.1. Rejection probabilities
When Assumption (A1) holds certain upper and lower bounds can be obtained for rejection probabilities with the help of Jensen’s inequality.
Lemma 29.
If Assumption (A1) holds, $\lambda\geq 1$ and $z_{1}\in\mathbb{S}$, then
$0$ $<w(z_{1})\frac{w(z_{1})+2\lfloor\lambda\rfloor-\lambda}{(w(z_{1})+\lfloor\lambda\rfloor)(w(z_{1})+\lfloor\lambda\rfloor-1)}\leq\epsilon(\lambda,z_{1})\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1},$
$\frac{1}{\lambda}$ $\leq b(\lambda)\leq\epsilon(\lambda)\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1},$
where
$b(\lambda)=\frac{1}{\lfloor\lambda\rfloor}-\frac{\lambda-\lfloor\lambda\rfloor}{(\lfloor\lambda\rfloor+1)\lfloor\lambda\rfloor}$
and for $\lambda>1$
$1<\frac{\lfloor\lambda\rfloor^{2}+3\lfloor\lambda\rfloor-\lambda+1}{\lfloor\lambda\rfloor^{2}-\lfloor\lambda\rfloor+\lambda-1}\leq\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)}\leq\frac{4\hat{w}+\lambda-1}{\lambda-1}.$
Proof.
Let $N\in\{2,3,\dots\}$, then
$\epsilon(N,z_{1})$ $=\int_{X^{N-1}}\frac{w(z_{1})}{\sum_{i=1}^{N}w(z_{i})}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)=1-\int_{X^{N-1}}\frac{\sum_{j=2}^{N}w(z_{j})}{\sum_{i=1}^{N}w(z_{i})}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
$=1-(N-1)\int_{X^{N-1}}\frac{w(z_{2})}{\sum_{i=1}^{N}w(z_{i})}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
$\leq 1-(N-1)\int_{X^{N-1}}\frac{w(z_{2})}{w(z_{1})+w(z_{2})+N-2}q(\mathrm{d}z_{2})$
$\leq 1-(N-1)\int_{X^{N-1}}\frac{w(z_{2})}{2\hat{w}+N-2}q(\mathrm{d}z_{2})$
$=1-\frac{N-1}{2\hat{w}+N-2}=\frac{2\hat{w}-1}{2\hat{w}+N-2},$
where the third equality follows from exchangeability of $z_{2},\dots,z_{N}$ under the integral and the first inequality from Jensen’s inequality. Therefore
$\epsilon(N)=\int_{\mathbb{S}}\epsilon(N,z_{1})\pi(\mathrm{d}z_{1})\leq\frac{2\hat{w}-1}{2\hat{w}+N-2},$
and for $N=1$ the inequality is clear for both cases. If $\lambda\geq 1$, then by Lemma 45 in Appendix D
$\epsilon(\lambda,z_{1})\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1},\qquad\epsilon(\lambda)\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1}.$
Let $N\in\mathbb{N}_{+}$ and $z_{1}\in\mathbb{S}$, then $\epsilon(1)=\epsilon(1,z_{1})=1$ and for $N\geq 2$ it holds by Jensen’s inequality, that
$\epsilon(N,z_{1})$ $=\int_{X^{N-1}}\frac{w(z_{1})}{\sum_{i=1}^{N}w(z_{i})}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\geq\frac{w(z_{1})}{w(z_{1})+N-1}$
and
$\epsilon(N)$ $\geq\int_{\mathbb{S}}\frac{w(z_{1})}{w(z_{1})+N-1}\pi(\mathrm{d}z_{1})=\int_{\mathbb{X}}\frac{w(z_{1})^{2}}{w(z_{1})+N-1}q(\mathrm{d}z_{1})\geq\frac{1}{N},$
where the last inequality follows from Jensen’s inequality, since $g:\mathbb{R}\to\mathbb{R}$,
$g(x)=\begin{cases}\frac{x^{2}}{x+N-1},&x>0\\ 0,&x\leq 0\end{cases}$
is convex. Thus, for $\lambda\geq 1$ it holds by Lemma 46 in Appendix D
$\epsilon(\lambda,z_{1})\geq w(z_{1})\frac{w(z_{1})+2\lfloor\lambda\rfloor-\lambda}{(w(z_{1})+\lfloor\lambda\rfloor)(w(z_{1})+\lfloor\lambda\rfloor-1)}>0$
and
$\epsilon(\lambda)\geq\frac{2\lfloor\lambda\rfloor+1-\lambda}{(\lfloor\lambda\rfloor+1)\lfloor\lambda\rfloor}=b(\lambda)\geq\frac{1}{\lambda}.$
Therefore, for $\lambda>1$
$1<\frac{\lfloor\lambda\rfloor^{2}+3\lfloor\lambda\rfloor-\lambda+1}{\lfloor\lambda\rfloor^{2}-\lfloor\lambda\rfloor+\lambda-1}=\frac{1+\frac{2\lfloor\lambda\rfloor+1-\lambda}{(\lfloor\lambda\rfloor+1)\lfloor\lambda\rfloor}}{1-\frac{2\lfloor\lambda\rfloor+1-\lambda}{(\lfloor\lambda\rfloor+1)\lfloor\lambda\rfloor}}\leq\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)}\leq\frac{1+\frac{2\hat{w}}{2\hat{w}+\lambda-1}}{1-\frac{2\hat{w}}{2\hat{w}+\lambda-1}}=\frac{4\hat{w}+\lambda-1}{\lambda-1}.\qed$
We have by Jensen’s inequality
(11) $\begin{split}\epsilon(\lambda)^{2}&\leq\int_{\mathbb{S}\times\mathbb{X}^{\lfloor\lambda\rfloor}}\bigg(\frac{\beta(\lambda)w(z_{1})}{\sum_{i=1}^{\lfloor\lambda\rfloor}w(z_{i})}+\frac{(1-\beta(\lambda))w(z_{1})}{\sum_{j=1}^{\lfloor\lambda\rfloor+1}w(z_{j})}\bigg)^{2}\bigg(\prod_{n=2}^{\lfloor\lambda\rfloor+1}q(\mathrm{d}z_{n})\bigg)\pi(\mathrm{d}z_{1}),\end{split}$
where $\beta(\lambda)=\lfloor\lambda\rfloor+1-\lambda\in(0,1]$. We shall define the function on the right-hand side of (11), which was already mentioned in Section 5: For Algorithm 2 the upper bound function of squared average rejection probability obtained by Jensen’s inequality is denoted by $\epsilon_{s}:[1,\infty)\to\mathbb{R}$,
(12) $\begin{split}\epsilon_{s}(\lambda)&=\int_{\mathbb{S}\times\mathbb{X}^{\lfloor\lambda\rfloor}}\bigg(\frac{\beta(\lambda)w(z_{1})}{\sum_{i=1}^{\lfloor\lambda\rfloor}w(z_{i})}+\frac{(1-\beta(\lambda))w(z_{1})}{\sum_{j=1}^{\lfloor\lambda\rfloor+1}w(z_{j})}\bigg)^{2}\bigg(\prod_{n=2}^{\lfloor\lambda\rfloor+1}q(\mathrm{d}z_{n})\bigg)\pi(\mathrm{d}z_{1})\\ &=\int_{\mathbb{W}_{\lfloor\lambda\rfloor+1}}w(z_{1})^{3}\bigg(\frac{\beta(\lambda)w(z_{\lfloor\lambda\rfloor+1})+\sum_{k=1}^{\lfloor\lambda\rfloor}w(z_{k})}{(\sum_{i=1}^{\lfloor\lambda\rfloor}w(z_{i}))(\sum_{j=1}^{\lfloor\lambda\rfloor+1}w(z_{j}))}\bigg)^{2}\bigg(\prod_{n=1}^{\lfloor\lambda\rfloor+1}q(\mathrm{d}z_{n})\bigg).\end{split}$
Lemma 30.
It holds that $\epsilon_{s}:[1,\infty)\to\mathbb{R}$ is strictly decreasing positive function, $\epsilon(\lambda)^{2}\leq\epsilon_{s}(\lambda)$ and $\epsilon_{s}(1)=1$. If Assumption (A1) holds, then $\epsilon_{s}(\lambda)\leq\lfloor\lambda\rfloor^{-1}\hat{w}$.
Proof.
First inequality follows from (11) and $\epsilon_{s}(1)=1$ by the definition. When $N\in\mathbb{N}_{+}$ and $N\leq\lambda_{1}<\lambda_{2}<N+1$, then by the definition of $\epsilon_{s}$, we have $\epsilon_{s}(\lambda_{1})>\epsilon_{s}(\lambda_{2})$. If $\lambda\in[N,N+1)$, then
$\epsilon_{s}(\lambda)-\epsilon_{s}(N+1)$
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\bigg(\frac{\beta(\lambda)w(z_{N+1})+\sum_{k=1}^{N}w(z_{k})}{(\sum_{j=1}^{N}w(z_{j}))(\sum_{l=1}^{N+1}w(z_{l}))}\bigg)^{2}-w(z_{1})^{3}\bigg(\frac{1}{\sum_{l=1}^{N+1}w(z_{l})}\bigg)^{2}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\frac{(\beta(\lambda)w(z_{N+1})+\sum_{k=1}^{N}w(z_{k}))^{2}-(\sum_{i=1}^{N}w(z_{i}))^{2}}{(\sum_{j=1}^{N}w(z_{j}))^{2}(\sum_{l=1}^{N+1}w(z_{l}))^{2}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\frac{\beta(\lambda)w(z_{N+1})[\beta(\lambda)w(z_{N+1})+2\sum_{k=1}^{N}w(z_{k})]}{(\sum_{j=1}^{N}w(z_{j}))^{2}(\sum_{l=1}^{N+1}w(z_{l}))^{2}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)>0.$
Therefore, $\epsilon_{s}$ is strictly decreasing. When Assumption (A1) holds, we have
$\epsilon_{s}(N)$ $=\int_{\mathbb{W}_{N}}w(z_{1})^{3}\bigg(\frac{1}{\sum_{i=1}^{N}w(z_{i})}\bigg)^{2}\bigg(\prod_{n=1}^{N}q(\mathrm{d}z_{n})\bigg)$
$\leq\hat{w}\int_{\mathbb{W}_{N}}\frac{w(z_{1})^{2}}{(\sum_{i=1}^{N}w(z_{i}))^{2}}\bigg(\prod_{n=1}^{N}q(\mathrm{d}z_{n})\bigg)\leq\hat{w}\int_{\mathbb{W}_{N}}\frac{w(z_{1})^{2}}{\sum_{i=1}^{N}w(z_{i})^{2}}\bigg(\prod_{n=1}^{N}q(\mathrm{d}z_{n})\bigg)$
$=\frac{\hat{w}}{N}\int_{\mathbb{W}_{N}}\frac{\sum_{j=1}^{N}w(z_{j})^{2}}{\sum_{i=1}^{N}w(z_{i})^{2}}\bigg(\prod_{n=1}^{N}q(\mathrm{d}z_{n})\bigg)\leq\frac{\hat{w}}{N},$
where the last equality follows from exchangeability of $z_{1},\dots z_{N}$ under the integral. Since $\epsilon_{s}$ is decreasing $\epsilon_{s}(\lambda)\leq\lfloor\lambda\rfloor^{-1}\hat{w}$. ∎
Let us define for Algorithm 2 the derivative of average rejection probability. The derivative function $\epsilon^{\prime}$ is undefined for $\lambda\in\mathbb{N}_{+}$, which can be seen from (10) and Lemma 18. We will use the same notation for the extended derivative function:
$\epsilon^{\prime}(\lambda)=\epsilon(\lfloor N\rfloor+1)-\epsilon(\lfloor N\rfloor),\qquad\lambda\in[1,\infty),$
which is a càdlàg function. The following bounds hold for this function:
Lemma 31.
The extended derivative function $\epsilon^{\prime}(\lambda):[1,\infty)\to\mathbb{R}$ is increasing negative function. Let $N\in\mathbb{N}_{+}$, then $\epsilon^{\prime}(x)=\epsilon^{\prime}(y)$ if $N\leq x\leq y<N+1$, and $\epsilon^{\prime}(x)<\epsilon^{\prime}(y)$ if $N\leq x<N+1\leq y$. If Assumption (A1) holds, then
$\frac{C}{(2\hat{w}+\lfloor\lambda\rfloor-1)^{2}}\leq|\epsilon^{\prime}(\lambda)|\leq\frac{\hat{w}}{\lfloor\lambda\rfloor(\lfloor\lambda\rfloor+1)},$
where $C=\int_{\mathbb{S}}w(x)\pi(\mathrm{d}x)\in[1,\hat{w}$].
Proof.
Let us first assume Assumption (A1) holds. Let $N\in\mathbb{N}_{+}$ and let us denote $B_{n}=\sum_{i=1}^{n}w(z_{i})$, then
$\epsilon(N)-\epsilon(N+1)$
$\quad=\int_{\mathbb{S}}\int_{\mathbb{X}^{N}}\frac{w(z_{1})w(z_{N+1})}{B_{N}B_{N+1}}\bigg(\prod_{n=2}^{N+1}q(\mathrm{d}z_{n})\bigg)\pi(\mathrm{d}z_{1})=\int_{\mathbb{W}_{N}\times\mathbb{X}}\frac{w(z_{1})^{2}w(z_{N+1})}{B_{N}B_{N+1}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$\quad\leq\hat{w}\int_{\mathbb{W}_{N}\times\mathbb{X}}\frac{w(z_{1})w(z_{N+1})}{B_{N}B_{N+1}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)=\frac{\hat{w}}{N}\int_{\mathbb{W}_{N}\times\mathbb{X}}\frac{B_{N}w(z_{N+1})}{B_{N}B_{N+1}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$\quad=\frac{\hat{w}}{N}\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{w(z_{N+1})}{B_{N+1}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)=\frac{\hat{w}}{N(N+1)}\int_{\mathbb{W}_{N+1}\times\mathbb{X}}\frac{B_{N+1}}{B_{N+1}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$\quad\leq\frac{\hat{w}}{N(N+1)},$
where the third equality follows from exchangeability of $z_{1},\dots,z_{N}$ and the last from exchangeability of $z_{1},\dots,z_{N+1}$ under the integral. We have also
$\epsilon(N,z_{1})-\epsilon(N+1,z_{1})$
$\quad=\int_{\mathbb{X}^{N}}\frac{w(z_{1})w(z_{N+1})}{B_{N}B_{N+1}}\bigg(\prod_{n=2}^{N+1}q(\mathrm{d}z_{n})\bigg)\geq\int_{\mathbb{X}^{N}}\frac{w(z_{1})w(z_{N+1})}{(B_{N+1})^{2}}\bigg(\prod_{n=2}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$\quad\geq\int_{\mathbb{X}^{N}}\frac{w(z_{1})w(z_{N+1})}{(2\hat{w}+\sum_{i=2}^{N}w(z_{i}))^{2}}\bigg(\prod_{n=2}^{N+1}q(\mathrm{d}z_{n})\bigg)\geq\int_{\mathbb{X}}\frac{w(z_{1})w(z_{N+1})}{(2\hat{w}+N-1)^{2}}q(\mathrm{d}z_{N+1})$
$\quad=\frac{w(z_{1})}{(2\hat{w}+N-1)^{2}},$
where the last inequality follows from Jensen’s inequality. Let us define $\nu=\mu|_{\mathbb{S}}$, then by Hölder’s inequality
$\Big(\int_{\mathbb{S}}\Big|\frac{\pi(x)}{\sqrt{q(x)}}\sqrt{q(x)}\Big|\nu(\mathrm{d}x)\Big)^{2}\leq\int_{\mathbb{S}}\Big|\frac{\pi(x)}{\sqrt{q(x)}}\Big|^{2}\nu(\mathrm{d}x)\int_{\mathbb{S}}\Big|\sqrt{q(y)}\Big|^{2}\nu(\mathrm{d}y),$
hence
$\int_{\mathbb{S}}w(x)\pi(\mathrm{d}x)=\int_{\mathbb{S}}\frac{\pi(x)^{2}}{q(x)}\nu(\mathrm{d}x)\geq\frac{\Big(\int_{\mathbb{S}}\pi(x)\nu(\mathrm{d}x)\Big)^{2}}{\int_{\mathbb{S}}q(x)\nu(\mathrm{d}x)}=\frac{1}{\int_{\mathbb{S}}q(x)\mu(\mathrm{d}x)}\geq 1.$
Rest of the claims follow from Lemma 18 and the definition of $\epsilon^{\prime}$. ∎
By summing up Lemma 31, we have that the extended derivative function $\epsilon^{\prime}:[1,\infty)\to\mathbb{R}$ is negative step function with positive steps on each integer and $\epsilon^{\prime}(\lambda)=O(\lambda^{-2})$. We can also give bounds for the holding probability:
Lemma 32.
If $(\mathbb{S})>1$, then for all $x\in\mathbb{S}$ and $\lambda\geq 1$
$\pi(\{x\})$ $<P_{\lambda}(x,\{x\})\leq\pi(\{x\})+(1-\pi(\{x\}))\frac{2\hat{w}}{2\hat{w}+\lambda-1},$
$\int_{\mathbb{S}}\pi(\{x\})\pi(\mathrm{d}x)$ $<\int_{\mathbb{S}}P_{\lambda}(x,\{x\})\pi(\mathrm{d}x)\leq\int_{\mathbb{S}}\pi(\{x\})\pi(\mathrm{d}x)+\Big(1-\int_{\mathbb{S}}\pi(\{x\})\pi(\mathrm{d}x)\Big)\frac{2\hat{w}}{2\hat{w}+\lambda-1}$
and for $\psi:[1,\infty)\to\mathbb{R}$,
$\psi(\lambda)=\frac{\int_{\mathbb{S}}[P_{\lambda}(x,\{x\})-\pi(\{x\})]\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}}\pi(\{x\})\pi(\mathrm{d}x)}$
holds
$0<\psi(\lambda)\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1},\qquad 1<\frac{1+\psi(\lambda)}{1-\psi(\lambda)}\leq\frac{4\hat{w}+\lambda-1}{\lambda-1},$
where all the upper bounds require Assumption (A1) to hold.
Proof.
We have $P_{\lambda}(x,\{x\})>\pi(\{x\})$ by Lemma 19 and Theorem 23. Rest of the inequalities follow from each other and Lemma 49 in Appendix E when $A=\{x\}$. ∎
8.2. Asymptotic variance
When Assumption (A1) holds, we can give upper and lower bounds for the asymptotic variance:
Lemma 33.
If Assumption (A1) holds, $\lambda>1$ and $f\in L_{0,+}^{2}(\pi)$, then
$\mathrm{var}_{\pi}(f)<\mathrm{var}(P_{\lambda},f)\leq\frac{4\hat{w}+\lambda-1}{\lambda-1}\mathrm{var}_{\pi}(f).$
Proof.
The transition probability $P_{\lambda}$ is $\pi$-reversible by Lemma 26, hence $\langle f|P_{\lambda}g\rangle_{\pi}=\langle P_{\lambda}f|g\rangle_{\pi}$ for all $g\in L^{2}(\pi)$, it also holds that $\langle g|P_{\lambda}g\rangle_{\pi}\geq 0$ and $\langle f|P_{\lambda}f\rangle_{\pi}>0$ by Theorem 20. Therefore
$\langle f|(I-P_{\lambda})^{-1}f\rangle_{\pi}=\sum_{k=0}^{\infty}\langle f|P_{\lambda}^{k}f\rangle_{\pi}=\sum_{j=0}^{\infty}\langle P_{\lambda}^{j}f|P_{\lambda}^{j}f\rangle_{\pi}+\sum_{k=0}^{\infty}\langle P_{\lambda}^{k}f|P_{\lambda}P_{\lambda}^{k}f\rangle_{\pi}>\langle f|f\rangle_{\pi}$
and thus
$\mathrm{var}(P_{\lambda},f)=2\langle f|(I-P_{\lambda})^{-1}f\rangle_{\pi}-\langle f|f\rangle_{\pi}>\langle f|f\rangle_{\pi}=\mathrm{var}_{\pi}(f).$
The upper bound can be found by combining [andrieu-lee-vihola, Proposition 32 in the supplement] with Lemma 49 in Appendix E. ∎
9. Approximate i-SIR in the case of atoms
We will show how approximation of the i-SIR kernel (Definition 5 in Section 4) is derived in the general case. Let us denote by $\hat{P}_{c}$ the transition probability in Lemma 37 in Appendix B, where $c$ is the holding probability (rejection probability when $\pi$ has no atoms), then it also holds that
(13) $\int\hat{P}_{c}(x,\{x\})\pi(\mathrm{d}x)=\int c+(1-c)\pi(\{x\})\pi(\mathrm{d}x).$
Finding the approximate kernel is done by solving $c$ from (13), where $\hat{P}_{c}$ is replaced with i-SIR transition probability $P_{\lambda}$. Then the single proposal Markov chain $\hat{P}_{c}$ with the corresponding holding probability $c$ is used as the approximate. The solution $c$ to (13) with the replacement is
(14) $\begin{split}\psi(\lambda)&=\frac{\int_{\mathbb{S}}[P_{\lambda}(x,\{x\})-\pi(\{x\})]\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}}\pi(\{x\})\pi(\mathrm{d}x)}\\ &=\frac{\int_{\mathbb{S}_{0}}\epsilon(\lambda,x)\pi(\mathrm{d}x)+\int_{\mathbb{S}_{+}}[P_{\lambda}(x,\{x\})-\pi(\{x\})]\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}.\end{split}$
Thus, we obtain the following:
Definition 34.
The approximate i-SIR transition $\tilde{Q}_{\lambda}$ of $P_{\lambda}$ is defined as
$\tilde{Q}_{\lambda}(x,A)=\big(1-\psi(\lambda)\big)\pi(A)+\psi(\lambda)\delta_{x}(A),$
where $\psi$ is given in (14).
Proposition 35.
If (A1) holds, then for any $f\in L^{2}(\pi)$ such that $\mathrm{var}_{\pi}(f)>0$
  1. (i)
    $\mathrm{var}(\tilde{Q}_{\lambda},f)=\frac{1+\psi(\lambda)}{1-\psi(\lambda)}\mathrm{var}_{\pi}(f).$
  2. (ii)
    $\mathrm{var}_{\pi}(f)\leq\mathrm{var}(\tilde{Q}_{\lambda},f)\leq\frac{4\hat{w}+\lambda-1}{\lambda-1}\mathrm{var}_{\pi}(f)$.
  3. (iii)
    $\lambda\mapsto\mathrm{var}(\tilde{Q}_{\lambda},f)$ is strictly convex and decreasing.
Proof.
Case (i) follows from Lemma 37 in Appendix B, case (ii) from Lemma 32 and case (iii) from Lemma 25. ∎
We note that Definition 5 and Proposition 6 in Section 4 correspond with Definition 34 and Proposition 35 when Assumption (A2) holds.
We use the following shorthand notation
(15) $V_{f}(\lambda)=\mathrm{var}(P_{\lambda},f),\qquad G_{f}(\lambda)=\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)}\mathrm{var}_{\pi}(f),\qquad H_{f}(\lambda)=\frac{1+\psi(\lambda)}{1-\psi(\lambda)}\mathrm{var}_{\pi}(f).$
Both functions $G_{f}$ and $H_{f}$ are well defined (Lemma 24 and 25) and have similar properties compared to the asymptotic variance $V_{f}$ when (A1) holds, namely they are strictly convex decreasing functions and they all behave between the same bounds (Theorem 4, Proposition 6 and 35). These imply that to some extent we can use $G_{f}$ and $H_{f}$ as approximates of $V_{f}$. Nevertheless, the smaller $\hat{w}$ gets, the closer $G_{f}$ and $H_{f}$ are to the asymptotic variance, and when $\hat{w}=1$, they are the same:
Proposition 36.
If $\pi=q$ almost surely, then the following holds for all $x\in\mathbb{S}$ and $\lambda\geq 1$
$P_{\lambda}(x,A)$ $=b_{\lambda}\mathbf{1}\left(x\in A\right)+(1-b_{\lambda})\pi(A),\qquad\epsilon(\lambda,x)=b_{\lambda},\qquad\epsilon(\lambda)=b_{\lambda},$
where $b_{\lambda}=\frac{1}{\lfloor\lambda\rfloor}-\frac{\lambda-\lfloor\lambda\rfloor}{(\lfloor\lambda\rfloor+1)\lfloor\lambda\rfloor}$ and for $f\in L_{0,+}^{2}(\pi)$
$\mathrm{var}(P_{\lambda},f)=G_{f}(\lambda)=H_{f}(\lambda)=\frac{1+b_{\lambda}}{1-b_{\lambda}}\mathrm{var}_{\pi}(f).$
Proof.
Since $\pi=q$ almost surely, we have $w=1$ almost surely in $\mathbb{S}$. Let $N\in\mathbb{N}_{+}$, then
$P_{N}(x,A)=\frac{1}{N}\mathbf{1}\left(x\in A\right)+\frac{N-1}{N}\pi(A),\qquad\epsilon(N,x)$ $=\frac{1}{N},\qquad\epsilon(N)=\frac{1}{N}.$
It holds for $\lambda\geq 1$ and $\beta(\lambda)=\lfloor\lambda\rfloor+1-\lambda$ that
$\beta(\lambda)\frac{1}{\lfloor\lambda\rfloor}+\big(1-\beta(\lambda)\big)\frac{1}{\lfloor\lambda\rfloor+1}=b_{\lambda},\qquad\beta(\lambda)\frac{\lfloor\lambda\rfloor-1}{\lfloor\lambda\rfloor}+\big(1-\beta(\lambda)\big)\frac{\lfloor\lambda\rfloor}{\lfloor\lambda\rfloor+1}=1-b_{\lambda}.$
Combining these with the definitions and Lemma 37 in Appendix B, we obtain the claim. ∎
Namely if $\pi=q$ almost surely, then Algorithm 2 gives the Markov chain in Lemma 37 in Appendix B.
10. Experiments
We investigate three questions in our experiments: the quality of the approximation of the asymptotic variance, the minimisation based on it, and the behaviour of the adaptive i-SIR algorithm in practice. Section 10.1 focuses on the two first questions, investigating the behaviour of i-SIR in finite state spaces $\mathbb{S}\subset\mathbb{R}$, where we can either calculate numerically or estimate efficiently the asymptotic variance of the i-SIR from its transition probability. Sections 10.2 and 10.3 investigate the use of the adaptive i-SIR in $\mathbb{S}=\mathbb{R}^{d}$: Section 10.2 presents a simple mixture example and Section 10.3 an example on Bayesian logistic regression.
10.1. Quality of asymptotic variance approximations in a finite state space
We first compare the behaviour of the asymptotic variance $V_{f}$ of the i-SIR algorithm with respect to the approximate asymptotic variances $G_{f}$ and $H_{f}$ given in (15), the corresponding loss functions and their minimisers. We consider different cost functions of type $c(\lambda)=a+\lambda$ for $a>0$, different target and proposal distributions and test functions.
In a finite state space (A1) always holds, and the asymptotic variance of a Markov chain can be calculated in principle using the spectral decomposition of its transition probability (Lemma 39 in Appendix B). In the case of i-SIR, the transition and rejection probabilities are not directly available, and direct calculation is too costly for large state space (and/or large number of proposals). Therefore, we consider two cases:
  1. (i)
    Direct calculations ($\mathbb{S}$ small; Experiments $10.1$).
  2. (ii)
    Monte Carlo approximations ($\mathbb{S}$ large; Experiment 10.1; see Proposition 55 in Appendix G).
We let $\lambda_{f}=\arg\min_{\lambda>1}c(\lambda)V_{f}(\lambda)$, and similarly $\lambda_{G_{f}}$ and $\lambda_{H_{f}}$ stand for the minimisers of the loss and the approximate loss functions using $G_{f}$ and $H_{f}$ in place of $V_{f}$, respectively. We define the following suboptimality factors:
$SO_{G}^{f}=\frac{V_{f}(\lambda_{G_{f}})}{V_{f}(\lambda_{f})}\qquad\text{and}\qquad SO_{H}^{f}=\frac{V_{f}(\lambda_{H_{f}})}{V_{f}(\lambda_{f})},$
which measure the effective loss of using the minimisers of the approximate losses instead of the true loss minimiser $\lambda_{f}$. For any (non-constant) $f\in L^{2}(\pi)$, we define its standardised version $\hat{f}(x)={\mathrm{var}^{-1/2}_{\pi}(f)}\big(f(x)-\pi(f)\big)$. Note that the minimisers $\lambda_{f}$, $\lambda_{G}$ and $\lambda_{H}$ are unaffected by the standardisation, and that $H_{\hat{f}}$ and $G_{\hat{f}}$ do not depend on the test function at all. We denote $\hat{G}=G_{\hat{f}}$, $\hat{H}=H_{\hat{f}}$, $\lambda_{\hat{G}}=\lambda_{G_{\hat{f}}}$ and $\lambda_{\hat{H}}=\lambda_{H_{\hat{f}}}$.
If the number of states $n=\#(\mathbb{S})=2$, then $H_{f}=V_{f}$ by Proposition 56 in Appendix G, but for $n\geq 3$, $V_{f}$ generally differs from $G_{f}$ and $H_{f}$. If the weight upper bound $\hat{w}$ is close to one, then Proposition 36 and the upper and lower bounds in Section 8 imply that all these functions are close to each other. But for large $\hat{w}$, these functions can differ more.
It is relatively easy to see from Algorithm 2 that the i-SIR chain can get stuck to states $s\in\mathbb{S}$ for which $w(s)$ is large. We denote by $Y_{u}^{M}=\{s\in\mathbb{S}:w(s)\geq M\}$ and $Y_{l}^{m}=\{s\in\mathbb{S}:w(s)\leq m\}$ the superlevel and sublevel sets of $w$, and consider the standardised versions of the following test functions: the identity function $f(x)=x$, the inverse weight function $g(x)=w^{-1}(x)$, the function
$h(x)=\begin{cases}x,&x\notin Y_{u}^{M}\\ \sum_{z\notin Y_{u}^{M}}z\frac{\pi(z)}{1-\pi(Y_{u}^{M})},&x\in Y_{u}^{M},\end{cases}$
the superlevel indicator $k(x)=\mathbf{1}\left(x\in Y_{u}^{M}\right)$ and the sublevel indicator $l(x)=\mathbf{1}\left(x\in Y_{l}^{m}\right)$. The function $\hat{f}$ admits its smallest and largest values at the endpoints of $\mathbb{S}$, and for $\hat{g}$, these occur at the largest and smallest weight, respectively. Function $\hat{h}$ is similar to $\hat{f}$, but $\hat{h}(s)=0$ whenever $w(s)\geq M$. In all the experiments, we calculate the asymptotic variance, its upper and lower bounds (Section 8), and the approximations $\hat{G}$ and $\hat{H}$ for $\lambda\in[2,150]$ with a spacing of $0.01$. In Experiments $10.1$ we have omitted from the tables results related to $\hat{H}$, since in these examples $\lambda_{\hat{H}}$ was equal to $\lambda_{\hat{G}}$ with the tested $a$ values.
Experiment 1
We consider the state space $\mathbb{S}=\{1,2,3,4,5\}$ and target $\pi$ and proposal $q$ shown in Figure 1(a). We choose $M=\hat{w}$ and $m=\min_{s\in\mathbb{S}}w(s)$, hence $Y_{u}^{M}=\{4\}$ and $Y_{l}^{m}=\{3\}$. The results of the experiment are illustrated in Figure 1; note the logarithmic scale in 1(b) and 1(c). The weight upper bound is $\hat{w}\approx 1.76$, so the lower and upper bounds are not far off from the functions in Figure 1(b) and 1(c), although there are some differences for test functions $\hat{k}$ and $\hat{l}$, which are concentrated on the extreme weights. The suboptimality factors in Table 1 show that their values stay close to $1$ for all test functions, even when $a$ increases.
Figure 1. Experiment 10.1: 1(a) probability mass functions $\pi$ (red) and $q$ (black). 1(b) functions $V_{\hat{f}}$ (red), $V_{\hat{g}}$ (orange), $V_{\hat{h}}$ (brown), $V_{\hat{k}}$ (purple), $V_{\hat{l}}$ (pink), $\hat{G}$ (green dashed), $\hat{H}$ (blue dashed) with upper/lower bounds (black dotted). 1(c) loss i.e. the functions in 1(b) multiplied by the cost $c(\lambda)=a+\lambda$.
a $\lambda_{\hat{G}}$ $\lambda_{\hat{f}}$ $SO_{G}^{\hat{f}}$ $\lambda_{\hat{g}}$ $SO_{G}^{\hat{g}}$ $\lambda_{\hat{h}}$ $SO_{G}^{\hat{h}}$ $\lambda_{\hat{k}}$ $SO_{G}^{\hat{k}}$ $\lambda_{\hat{l}}$ $SO_{G}^{\hat{l}}$
0 3 3 1 3 1 3 1 3 1 2 1.06
0.1 3 3 1 3 1 3 1 3 1 2 1.04
1 3 3 1 3 1 3 1 4 1.02 3 1
2 4 4 1 4 1 4 1 4 1 3 1.04
5 5 5 1 5 1 5 1 6 1.00 4 1.03
10 6 6 1 6 1 6 1 7 1.01 5 1.01
20 8 8 1 8 1 8 1 9 1.01 6 1.01
Table 1. Experiment 10.1: Minimisers and their corresponding suboptimality factors.
Experiment 2
Next, we consider state space $\mathbb{S}=\{1,2,\ldots,6\}$ and the probability mass functions $\pi,q:\mathbb{S}\to[0,1]$ shown in Figure 2(a), where states $1$ to $4$ have the same mass as in Experiment 10.1. As in Experiment 10.1 we choose $M=\hat{w}$ and $m=\min_{s\in\mathbb{S}}w(s)$, hence $Y_{u}^{M}=\{6\}$ and $Y_{l}^{m}=\{5\}$. The weights can take much larger values $w(6)=\hat{w}\approx 430.99$ than in Experiment 10.1, so the lower and upper bounds are much more apart, and the shape of the asymptotic variance varies among the test functions. The approximations $\hat{H}$ and $\hat{G}$ are fairly good for $\hat{h}$ and $\hat{l}$, but have a noticeable discrepancy for $\hat{f}$, $\hat{g}$ and $\hat{k}$. This reflects also in the discrepancy of the minimisers, which is sometimes ten-fold, and the suboptimality factors shown in Table 2.
Figure 2. Experiment 10.1: 2(a) probability mass functions $\pi$ (red) and $q$ (black). 2(b) functions $V_{\hat{f}}$ (red), $V_{\hat{g}}$ (orange), $V_{\hat{h}}$ (brown), $V_{\hat{k}}$ (purple), $V_{\hat{l}}$ (pink), $\hat{G}$ (green dashed), $\hat{H}$ (blue dashed) with upper/lower bounds (black dotted). 2(c) loss i.e. the functions in 2(b) multiplied by the cost $c(\lambda)=a+\lambda$.
a $\lambda_{\hat{G}}$ $\lambda_{\hat{f}}$ $SO_{G}^{\hat{f}}$ $\lambda_{\hat{g}}$ $SO_{G}^{\hat{g}}$ $\lambda_{\hat{h}}$ $SO_{G}^{\hat{h}}$ $\lambda_{\hat{k}}$ $SO_{G}^{\hat{k}}$ $\lambda_{\hat{l}}$ $SO_{G}^{\hat{l}}$
0 3 21 1.37 19 1.35 3 1 33 1.41 6 1.13
0.1 3 22 1.40 20 1.39 3 1 34 1.45 6.15 1.15
1 3 29 1.75 26 1.73 3 1 45 1.84 8 1.32
2 4 35 1.70 32 1.68 4 1 55 1.80 9.7 1.24
5 5 49 2.00 44.85 1.96 5 1 78 2.16 13.1 1.31
10 6 66 2.37 60 2.31 6 1 104 2.63 17.48 1.39
20 8 91 2.68 82 2.59 8 1 143 3.07 23.82 1.39
Table 2. Experiment 10.1: Minimisers and their corresponding suboptimality factors.
Experiment 3
We consider again the same state space $\mathbb{S}=\{1,2,\ldots,6\}$. The probability mass functions $\pi,q:\mathbb{S}\to[0,1]$ shown in Figure 3(a) are the same as in Experiments 10.1 and 10.1 for states $1$ to $4$, and $\hat{w}$ is same as in Experiment 10.1, but $\pi(6)$ and $q(6)$ are $1000$ times smaller than in Experiment 10.1. We choose $M$ and $m$ as before, hence $Y_{u}^{M}=\{6\}$ and $Y_{l}^{m}=\{3\}$. The results in Figure 3 and Table 3 are similar to Experiment 10.1 for most test functions, but the results for $\hat{k}$ resemble those in Experiment 10.1: as $a$ increases, the suboptimality factor increases.
Figure 3. Experiment 10.1: 3(a) probability mass functions $\pi$ (red) and $q$ (black). 3(b) functions $V_{\hat{f}}$ (red), $V_{\hat{g}}$ (orange), $V_{\hat{h}}$ (brown), $V_{\hat{k}}$ (purple), $V_{\hat{l}}$ (pink), $\hat{G}$ (green dashed), $\hat{H}$ (blue dashed) with upper/lower bounds (black dotted). 3(c) loss i.e. the functions in 3(b) multiplied by the cost $c(\lambda)=a+\lambda$.
a $\lambda_{\hat{G}}$ $\lambda_{\hat{f}}$ $SO_{G}^{\hat{f}}$ $\lambda_{\hat{g}}$ $SO_{G}^{\hat{g}}$ $\lambda_{\hat{h}}$ $SO_{G}^{\hat{h}}$ $\lambda_{\hat{k}}$ $SO_{G}^{\hat{k}}$ $\lambda_{\hat{l}}$ $SO_{G}^{\hat{l}}$
0 3 3 1 3 1 3 1 30 1.41 2 1.06
0.1 3 3 1 3 1 3 1 32 1.45 2 1.04
1 3 3 1 3 1 3 1 43 1.82 3 1
2 4 4 1 4 1 4 1 52 1.79 3 1.04
5 5 5 1 5 1 5 1 73 2.14 4 1.03
10 6 6 1 6 1 6 1 98 2.60 5 1.01
20 8 9 1.00 9 1.00 8 1 136 3.02 6 1.01
Table 3. Experiment 10.1: Minimisers and their corresponding suboptimality factors.
Experiment 4
Here, we let $\mathbb{S}=\{1,2,3,4,5\}$ and consider the probability mass functions $\pi,q:\mathbb{S}\to[0,1]$ shown in Figure 4(a). The proposal $q$ is the same as in Experiment 10.1 but $\pi$ is modified: we divide the mass from state $5$ evenly to the other states in a way that $w(5)=1/d$, where $d$ is $\hat{w}$ of Experiment 10.1 and 10.1. The weights are small $w(4)=\hat{w}\approx 2.50$. We choose $M$ and $m$ as before, hence $Y_{u}^{M}=\{4\}$ and $Y_{l}^{m}=\{5\}$. With test functions $\hat{g}$ and $\hat{l}$, which take their extremal values at state 5, the results differ from earlier experiments: with large $a$, the cost function minimisers are much smaller than with approximations. These functions admit their extreme values at $5$, which the i-SIR chain does not visit often. For the rest of the test functions, the losses looks similar to the approximate losses as the results in Table 4 confirm.
Figure 4. Experiment 10.1: 4(a) probability mass functions $\pi$ (red) and $q$ (black). 4(b) functions $V_{\hat{f}}$ (red), $V_{\hat{g}}$ (orange), $V_{\hat{h}}$ (brown), $V_{\hat{k}}$ (purple), $V_{\hat{l}}$ (pink dashed), $\hat{G}$ (green dashed), $\hat{H}$ (blue dashed) with upper/lower bounds (black dotted). 4(c) loss i.e. the functions in 4(b) multiplied by the cost $c(\lambda)=a+\lambda$.
a $\lambda_{\hat{G}}$ $\lambda_{\hat{f}}$ $SO_{G}^{\hat{f}}$ $\lambda_{\hat{g}}$ $SO_{G}^{\hat{g}}$ $\lambda_{\hat{h}}$ $SO_{G}^{\hat{h}}$ $\lambda_{\hat{k}}$ $SO_{G}^{\hat{k}}$ $\lambda_{\hat{l}}$ $SO_{G}^{\hat{l}}$
0 3 4 1.01 3 1 3 1 4 1.01 3 1
0.1 3 4 1.02 3 1 3 1 4 1.02 3 1
1 4 4 1 3 1.08 4 1 5 1.01 3 1.08
2 5 5 1 3 1.15 5 1 5 1 3 1.15
5 6 7 1.00 4 1.14 6 1 7 1.01 4 1.15
10 8 8 1 4 1.19 7 1.00 9 1.00 4 1.19
20 10 11 1.00 5 1.17 9 1.00 11 1.00 5 1.17
Table 4. Experiment 10.1: Minimisers and their corresponding suboptimality factors.
Experiment 5
Contrary to the previous, purely artificial examples, this example is a discretisation of a continuous state model. We consider $\mathbb{S}$ which is a discretisation of [3,3][-3,3] to 61 equally spaced states, and $\pi$ and $q$ correspond to discretisation of two normal distributions $N(0,1/4)$ and $N(0,1)$, respectively, illustrated in Figure 5(a).
For integers $N\geq 2$ we have used Monte Carlo estimation to estimate the transition and rejection probabilities (i.e. (20) and (21) of Proposition 55 in Appendix G) by simulating $Z_{1}^{N},\dots,Z_{100,000}^{N}\sim\mathrm{Multinom}(N-1,Q)$. We calculate both $P_{N}$ and $\epsilon(N)$ using the same $(Z_{i}^{N})$, and also create new $Z_{i}^{N+1}$ conditional on $Z_{i}^{N}$. The estimates $P_{\lambda}$ and $\epsilon(\lambda)$ for non-integer $\lambda\geq 1$ are derived from these.
The weight upper bound $\hat{w}\approx 2.00$ is fairly small. We choose $Y_{u}^{M}$ and $Y_{l}^{m}$, where $M=1.9$ and $m=0.2$, for which $\pi(Y_{u}^{M})\approx 0.23$ and $\pi(Y_{l}^{m})\approx 0.012$. Figures 5(b) and 5(c) show that the asymptotic variances have some variation, but Table 5 indicates that all the suboptimality factors are relatively small. With $\hat{g}$, the factors are the highest, but as $a$ grows, they get smaller.
Figure 5. Experiment 5: 5(a) probability mass functions $\pi$ (red) and $q$ (black). 5(b) functions $V_{\hat{f}}$ (red), $V_{\hat{g}}$ (orange), $V_{\hat{h}}$ (brown dashed), $V_{\hat{k}}$ (purple), $V_{\hat{l}}$ (pink), $\hat{G}$ (green dashed), $\hat{H}$ (blue dashed) with upper/lower bounds (black dotted). 5(c) loss i.e. the functions in 5(b) multiplied by the cost $c(\lambda)=a+\lambda$.
a $\lambda_{\hat{G}}$ $\lambda_{\hat{f}}$ $SO_{G}^{\hat{f}}$ $\lambda_{\hat{g}}$ $SO_{G}^{\hat{g}}$ $\lambda_{\hat{h}}$ $SO_{G}^{\hat{h}}$ $\lambda_{\hat{k}}$ $SO_{G}^{\hat{k}}$ $\lambda_{\hat{l}}$ $SO_{G}^{\hat{l}}$
0 3 3 1 2 1.47 3 1 3 1 2 1.04
0 3 3 1 2 1.47 3 1 3 1 2 1.04
0.1 3 3 1 2 1.45 3 1 3 1 2 1.02
1 4 3 1.01 2 1.63 3 1.01 4 1 3 1.12
2 4 4 1 2 1.47 4 1 5 1.02 3 1.08
5 6 5 1.02 2 1.54 5 1.02 6 1 3 1.16
10 7 6 1.01 2 1.39 6 1.01 8 1.01 4 1.13
20 9 8 1.01 2 1.29 8 1.01 10 1.01 5 1.11
Table 5. Experiment 5: Minimisers and their corresponding suboptimality factors.
10.2. Mixture distribution
We run our adaptive i-SIR (Algorithm 3) on a 7-dimensional Gaussian mixture target example from [cardoso-samsonov-thin-moulines-olsson]. The modes of the target are $m_{1}=(1,\ldots,1)^{T}$ and $m_{2}=(-2,0,\ldots,0)^{T}$ and the components have identity covariance The proposal has zero mean and identity shape student’s $t$ distribution with $\nu=3$ degrees of freedom. We consider the same two test functions as in [cardoso-samsonov-thin-moulines-olsson]: $f_{1}(x)=x_{1}$ and $f_{2}(x)=\mathbf{1}\left(x\in A\right)-\mathbf{1}\left(x\in B\right)$, with sets $A=[-2,6]\times[-1,1]^{6}$ and $B=[0.75,1.25]\times[1,2]\times[-0.1,0.1]^{5}$.
The methods are implemented in the Julia programming language [julia], using its native parallelisation to the i-SIR proposal/density evaluation. The number of threads is set to four, which is the number of performance cores in the Macbook Pro 2020 with Apple M1 processor, where the tests are run. We first do pilot runs of $n=10,000$ iterations with i-SIR having fixed $N_{i}=2^{i}+1$ for $i\in\{2,\ldots,13\}$. The runs are timed, and a linear regression model is fit to the times $T_{i}=a+bN_{i}+\epsilon_{i}$, where $\epsilon_{i}$ are the residuals (Figure 6). The cost function is formed by scaling the fitted model: $c(\lambda)=(a/b)+\lambda$. After this, we run the adaptive i-SIR for the model. Figure 6 (middle) shows the evolution of the adapted $\lambda_{k}$ and the trace plot of the first coordinate of the samples, after a 10% burn-in.
Figure 6. Average running time $T_{i}$ for different $N_{i}$ and the fitted regression line (left), the evolution of $\lambda_{k}$ (middle) and the trace plot of the first coordinate of adaptive i-SIR output (right) in the mixture example of Section 10.2.
Then, we run i-SIRs with $N$ fixed to each $N_{i}$ and the final value of the adaptation for one million iterations. We use the initial sequence estimator to estimate the asymptotic variance [geyer-practical]. Figure 7 (left) shows the integrated autocorrelation time (IACT) estimates for the test functions and based on our approximation. The IACT of $f_{2}$ behaves close to the approximation, whereas the IACT of $f_{1}$ becomes close to the approximation only for higher number of proposals $N$. Figure 7 (right) the estimated inverse relative efficiencies [glynn-whitt], that is, average running time per iteration times asymptotic variance. Note that both the $x$ and $y$ axes are logarithmic. The vertical line indicates the value of the number of proposals corresponding to the final value of adaptation. The adaptation minimises the approximate IRE, and approximately minimises the IRE of the second test function. For the first test function, a greater number of proposals would be optimal. However, the loss of efficiency is not large.
Figure 7. The estimated integrated autocorrelation times (left) and the inverse relative efficiencies and the loss (right) in the mixture example of Section 10.2. The vertical grey line indicates the terminal value of adaptation.
10.3. Logistic regression
Our final example is a Bayesian logistic regression model on the Wisconsin Diagnostic Breast Cancer dataset (doi:10.24432/C5DW2B, [street-wolberg-mangasarian]), consisting of 569 patients’ data, with a binary outcome variable (benign/malignant), and with 30 covariates plus intercept. We use a Gaussian prior $\mathrm{pr}(x)=N(x;0,20I_{31})$ for the regression coefficients. We first find the Laplace approximation $\hat{\pi}(x)=N(x;\mu,\Sigma)$ of the posterior, and use a (‘defensive’) mixture $q(x)=0.1\mathrm{pr}(x)+0.9\hat{\pi}(x)$ as the proposal; this guarantees bounded weights $w(x)=\pi(x)/q(x)$ as long as the likelihood is bounded. The test functions are posterior density values $f_{1}(x)=\pi(x)$ and $f_{2}(x)=\|x-\mu\|$.
We use the same implementation as in Section 10.2, the same number of preliminary samples, but the actual simulations are now ten million iterations long. Figure 8 illustrates the results similar to Figure 6 (middle) and 7. Also here, the adaptive algorithm appears to minimise the approximate IRE, and the IACT and IRE of the first test function follows the same shape as the approximation. The IRE of $f_{2}$ seems to be roughly minimised, but because there remains fluctuations in the estimated IACT of $f_{2}$, no clear conclusions can be drawn.
Figure 8. The evolution of $\lambda_{k}$ (left), the estimated integrated autocorrelation times (middle) and the inverse relative efficiencies and the loss (right) in the logistic regression example of Section 10.3. The vertical grey line indicates the terminal value of adaptation.
11. Discussion
We presented an i-SIR algorithm with adaptive number of proposals, based on an acceptance rate based asymptotic variance approximation. The approximate asymptotic has similar properties (convexity and monotonicity) as the true i-SIR asymptotic variance, and our experiments suggest that it often leads to an appropriate choice of the number of proposals. We focused on the adaptation with an affine cost, which is assumed to be known. Our method can accommodate other forms of (smooth) cost functions, which might be appropriate for instance when run times have a large variation — because the i-SIR selection step requires to wait for the slowest sampling process to finish. We estimated the cost based on a pilot run with a range of number of proposals, by simply fitting a linear model to the observed running times. It might be possible to develop an online strategy to learn and/or refine the cost function based on observed run times.
We used an approximate asymptotic variance based on modelling i-SIR as ‘lazy perfect sampling.’ The approximation was derived for the non-atomic case (A2), but our experiments suggest that it is close to an approximation specifically tailored for discrete state spaces (cf. Section 9). While the approximation appeared to work well in many settings, it can be a bit crude, because sometimes the i-SIR chain can get stuck for longer time than anticipated by the i-SIR approximation. One approach to generalise our approximation might be to use a similar ‘slowed down’ perfect sampling, but in a semi-Markov manner: replacing the geometric holding times with some other, heavier tailed distribution. This could potentially account better for the excess ‘stickiness’ due to varying weights. Another possible approach is to investigate the combination of asymptotic variance upper bound based criteria [see cardoso-samsonov-thin-moulines-olsson], that is, to estimate the running mean of the weights to estimate the upper bound $W_{i}\leq\hat{w}$ use some combination of the bounds in the adaptation.
We presented our algorithm for the simplest possible variant of the i-SIR, and could be elaborated and extended in a number of ways. For instance, it is possible to use develop lower variance estimators of $\epsilon$ and $\epsilon^{\prime}$ by exchangeability of $Y_{k}^{2:\bar{N}_{k}j}$: the estimates (3) and (4) remain valid if the last item $Y_{k}^{\bar{N}_{k}}$ is changed to each of the proposals. It is possible to slightly increase the acceptance rate of the i-SIR by the forced-move variant [liu, chopinsingh2015]. We believe that our method could be used also with the forced-move extension, but we are unaware which theoretical properties remain valid.
Instead of using only the current value of the i-SIR Markov chain, it is possible to use the entire weighted sample in estimation; see for instance [cardoso-samsonov-thin-moulines-olsson]. It might be possible to modify our adaptation to this setting, but using all proposals leads to different efficiency, and this must be accounted for by developing a new asymptotic variance proxy. The i-SIR can be extended to non-independent proposals [mendes-scharth-kohn]; see also [karppinen-vihola], who also suggest covariance adaptation in this context. This setting also requires a different approximate asymptotic variance. Finally, adapting the proposal distribution, for instance with ideas from importance sampling [bugallo-elvira-martino-luengo-miguez-djuric], together with the number of proposals might be possible.
The approximate asymptotic variance which we use as a basis for adaptation leads to a loss function whose minimum does not depend on the test function $f$. We believe that this is often favourable (when we care about multiple test functions), but if we focus on one specific test function $f$, it may be possible to devise another function specific strategy for instance based on asymptotic correlations (the mean square jumping distance).
Acknowledgements
The authors were supported by the Research Council of Finland (grants 346311 & 364216), and wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.
Appendix A Details about the setup
In our setup we have a sigma-finite measure $\mu$ and probability measures $\Pi$ and $\mathbb{Q}$ on a measurable space $(\mathbb{X},\mathcal{X})$ such that $\Pi\ll\mathbb{Q}\ll\mu$. The Radon-Nikodym derivatives of $\Pi$ and $\mathbb{Q}$ with respect to $\mu$ shall be denoted by $\pi$ and $q$. These derivatives will be chosen so that $\pi,q\in[0,\infty)$ and $q(x)>0$ if $\pi(x)>0$, which can be done. The support of $\pi$ is denoted by $\mathbb{S}$. The Radon-Nikodym derivative of $\Pi$ with respect to $\mathbb{Q}$ will be denoted by $w$ i.e. $\pi=wq$ $\mu$-almost everywhere, which can be chosen to be
$w(x)=\begin{cases}\frac{\pi(x)}{q(x)},&x\in\mathbb{S}\\ 0&x\notin\mathbb{S}.\end{cases}$
For this weight function holds $w\in[0,\infty)$. We say that $x\in\mathbb{X}$ is an atom for measure $\mu$ when $\mu(\{x\})>0$. For $\mathbb{Q}$ this condition is equivalent to $q(x)>0$. Note that $\mathbb{Q}\ll\mu$ implies that if $x$ is not an atom for $\mu$, then it is neither for $\mathbb{Q}$. However an atom for $\mu$ doesn’t require to be an atom for $\mathbb{Q}$. Similar conditions also hold for $\Pi$ with respect to $\mathbb{Q}$. Note that in general $q(x)=\mathbb{Q}(\{x\})$ doesn’t hold for an atom of $\mathbb{Q}$, but for instance when $\mu$ restricted to the atoms of $\mathbb{Q}$ is counting measure, then $q(x)=\mathbb{Q}(\{x\})$. Especially when $\mu$ is Lebesgue measure/counting measure on $\mathbb{R}$ we can have the usual probability density functions/probability mass functions as the Radon-Nikodym derivatives of $\pi$ and $q$.
Appendix B Asymptotic variance results
Recall that the asymptotic variance of a Markov chain $P$ and function $f\in L_{0}^{2}(\pi)$ is
(16) $\mathrm{var}(P,f)=2\sum_{k=0}^{\infty}\langle f|P^{k}f\rangle_{\pi}-\|f\|_{\pi}^{2},$
whenever the sum on the right is well-defined [cf. tierney-note].
Lemma 37.
Let $\pi$ be a probability distribution on a general state space $E$, and let $\epsilon\in(0,1)$. Consider the Markov transition probability:
$P_{\epsilon}(x,A)=\epsilon\mathbf{1}\left(x\in A\right)+(1-\epsilon)\pi(A).$
The asymptotic variance of a function $f\in L_{0}^{2}(\pi)$ is:
$\mathrm{var}(P_{\epsilon},f)=\frac{1+\epsilon}{1-\epsilon}\|f\|_{\pi}^{2}$
Proof.
Let us have i.i.d. random variables $U_{1},\dots,U_{n}\sim Unif(0,1)$ and $X_{0},Z_{1},\dots,Z_{n}\sim\pi$ that are independent of each other. Let
$X_{n}=X_{n-1}I(U_{n}\leq\epsilon)+Z_{n}I(U_{n}>\epsilon)$
for $n\geq 1$, hence $X_{n}\sim\pi$. Let us standardise the function $f$ and denote $g(x)=\frac{f(x)}{\sqrt{\|f\|_{\pi}^{2}}}$. We have for all $n\geq 1$, that
$g(X_{n})=g(X_{n-1})\mathbf{1}\left(U_{n}\leq\epsilon\right)+g(Z_{n})\mathbf{1}\left(U_{n}>\epsilon\right),$
hence $g(X_{n})$ are identically distributed and it holds for all $m\geq 0$ that $\mathbb{E}[g(X_{m})]=0$ and $\mathbb{E}[g(X_{m})^{2}]=1$. By induction we get for $0\leq m<n$, that
$g(X_{n})=g(X_{m})\mathbf{1}\left(R_{m+1:n}\right)+\mathbf{1}\left(n-m\geq 2\right)\sum_{i=m+1}^{n-1}g(Z_{i})\mathbf{1}\left(A_{i}\cap R_{i+1:n}\right)+g(Z_{n})\mathbf{1}\left(A_{n}\right),$
where
$R_{j:k}=\{U_{j}\leq\epsilon,\dots,U_{k}\leq\epsilon\},\qquad A_{j}=\{U_{j}>\epsilon\}.$
Therefore
$\mathbb{E}[g(X_{m})g(X_{n})]$
$\quad=\mathbb{E}[g(X_{m})g(X_{m})\mathbf{1}\left(R_{m+1:n}\right)]+\mathbf{1}\left(n-m\geq 2\right)\sum_{i=m+1}^{n-1}\mathbb{E}[g(X_{m})g(Z_{i})\mathbf{1}\left(A_{i}\cap R_{i+1:n}\right)]$
$\qquad+\mathbb{E}[g(X_{m})g(Z_{n})\mathbf{1}\left(A_{n}\right)]$
$\quad=\mathbb{E}[g(X_{m})^{2}]\mathbb{E}[\mathbf{1}\left(R_{m+1:n}\right)]=\mathbb{E}[g(X_{m})^{2}]\epsilon^{n-m}=\epsilon^{n-m}$
and it follows that
$n\cdot\mathrm{var}\Big(\frac{\sum_{i=1}^{n}g(X_{i})}{n}\Big)$ $=\frac{1}{n}\mathbb{E}\Big[\Big(\sum_{i=1}^{n}g(X_{i})\Big)^{2}\Big]=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbb{E}[g(X_{i})g(X_{j})]$
$=\frac{1}{n}\Big(\sum_{i=1}^{n}\mathbb{E}[g(X_{i})^{2}]+2\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\mathbb{E}[g(X_{i})g(X_{j})]\Big)$
$=1+\frac{2}{n}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\epsilon^{j-i}=1+\frac{2}{n}\sum_{i=1}^{n}(n-i)\epsilon^{i}.$
We have $\sum_{i=1}^{\infty}\epsilon^{i}=\frac{\epsilon}{1-\epsilon}$, hence by Kronecker
$\sum_{i=1}^{n}\frac{i\epsilon^{i}}{n}\xrightarrow{n\to\infty}0.$
Thus
$\lim_{n\to\infty}n\cdot\mathrm{var}\Big(\frac{\sum_{i=1}^{n}g(X_{i})}{n}\Big)$ $=1+2\sum_{i=1}^{\infty}\epsilon^{i}=\frac{1+\epsilon}{1-\epsilon}$
and it follows that
$\lim_{n\to\infty}n\cdot\mathrm{var}\Big(\frac{\sum_{i=1}^{n}f(X_{i})}{n}\Big)$ $=\frac{1+\epsilon}{1-\epsilon}\|f\|_{\pi}^{2}.$
Since $(X_{n})_{n=0}^{\infty}$ is Markov chain with Markov transition probability $P_{\epsilon}$ and initial distribution $\pi$, we have
$\mathrm{var}(P_{\epsilon},f)=\frac{1+\epsilon}{1-\epsilon}\|f\|_{\pi}^{2}.\qed$
When the state space is finite, we have the following formulation:
Lemma 38.
Let $P=[P_{i,j}]_{i,j=1,\dots,n}$ be a transition probability matrix, such that $P$ is irreducible, aperiodic and $\pi$-reversible with respect to some probability vector $\pi\in\mathbb{R}^{n}$. Let $(X_{k})_{k=0}^{\infty}$ be a Markov chain in state space $\mathbb{S}=(s_{1},\dots,s_{n})$ with transition probability $P$ and initial distribution $\pi$, then it holds for $f:\mathbb{S}\to\mathbb{R}$ that
$\mathrm{var}(P,\bar{f})=\sum_{i\in\{1,\dots,n\}\setminus m}\langle u_{i}|\bar{f}\rangle_{\pi}^{2}\frac{1+\lambda_{i}}{1-\lambda_{i}},$
where $\lambda_{1},\dots,\lambda_{n}$ are eigenvalues and $u_{1},\dots,u_{n}$ eigenvectors of matrix $P$ with $\lambda_{m}=1$ for some $m\in\{1,\dots,n\}$, $|\lambda_{i}|<1$ for all $i\neq m$, $u_{m}=(1,\dots,1)$, where the vectors $u_{1},\dots,u_{n}$ form an orthonormal basis for $(\mathbb{R}^{n},\langle\cdot|\cdot\rangle_{\pi})$.
Proof.
We obtain the given eigenvalues and eigenvectors by [levin-peres-wilmer, Lemma 12.1, Lemma 12.2]. By these properties $P$ is diagonalizable and
$P^{k}=CD^{k}C^{-1},$
where $D=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ and $C=[u_{1},\dots,u_{n}]$. Let us denote
$\Pi=\mathrm{diag}(\pi_{1},\dots,\pi_{n}),$
then
$\Pi^{1/2}P^{k}\Pi^{-1/2}=\Pi^{1/2}CD^{k}C^{-1}\Pi^{-1/2}=QD^{k}Q^{-1},$
where $Q=\Pi^{1/2}C=[\Pi_{j}^{1/2}u_{ij}]_{i,j=1,\dots,n}$. We can easily verify $Q^{T}Q=I$ i.e. $Q^{T}=Q^{-1}$, since
$\langle u_{i}|u_{j}\rangle_{\pi}=\begin{cases}1&i=j\\ 0&i\neq j.\end{cases}$
We have
$\langle\bar{f}|P^{k}\bar{f}\rangle_{\pi}=\bar{f}^{T}\Pi P^{k}\bar{f}=\bar{f}^{T}\Pi^{1/2}QD^{k}Q^{-1}\Pi^{1/2}\bar{f}=\bar{f}^{T}\Pi^{1/2}QD^{k}Q^{T}\Pi^{1/2}\bar{f}=e^{T}D^{k}e,$
where $e=Q^{T}\Pi^{1/2}\bar{f}$ and thus
$e=Q^{T}\Pi^{1/2}\bar{f}=C^{T}\Pi^{1/2}\Pi^{1/2}\bar{f}=C^{T}\Pi\bar{f}=[\langle u_{i}|\bar{f}\rangle_{\pi}]_{i=1,\dots,n},$
i.e. $e_{i}=\langle u_{i}|\bar{f}\rangle_{\pi}$, where $e_{m}=0$, since $\bar{f}\in u_{m}^{\perp}$ and
$\|\bar{f}\|_{\pi}^{2}=\langle\bar{f}|\bar{f}\rangle_{\pi}=e^{T}e=\sum_{i=1}^{n}e_{i}^{2}=\sum_{i=1}^{n}\langle u_{i}|\bar{f}\rangle_{\pi}^{2}=\sum_{i\in\{1,\dots n\}\setminus m}\langle u_{i}|\bar{f}\rangle_{\pi}^{2}.$
Let us define diagonal matrix $R=\mathrm{diag}(\lambda_{1},\dots,\lambda_{m-1},0,\lambda_{m+1},\dots,\lambda_{n})$, then
$\sum_{k=0}^{\infty}R^{k}=\mathrm{diag}\Big(\frac{1}{1-\lambda_{1}},\dots,\frac{1}{1-\lambda_{m-1}},0,\frac{1}{1-\lambda_{m+1}},\dots,\frac{1}{1-\lambda_{n}}\Big),$
and
$\sum_{k=0}^{\infty}\langle\bar{f}|P^{k}\bar{f}\rangle=\sum_{k=0}^{\infty}e^{T}D^{k}e=\sum_{k=0}^{\infty}e^{T}R^{k}e=\sum_{i\in\{1,\dots,n\}\setminus m}\langle u_{i}|\bar{f}\rangle_{\pi}^{2}\frac{1}{1-\lambda_{i}}.$
Therefore
$\mathrm{var}(P,\bar{f})$ $=2\sum_{k=0}^{\infty}\langle\bar{f}|P^{k}\bar{f}\rangle_{\pi}-\|\bar{f}\|_{\pi}^{2}=\sum_{i\in\{1,\dots,n\}\setminus m}\langle u_{i}|\bar{f}\rangle_{\pi}^{2}\Big[\frac{2}{1-\lambda_{i}}-1\Big]$
$=\sum_{i\in\{1,\dots,n\}\setminus m}\langle u_{i}|\bar{f}\rangle_{\pi}^{2}\frac{1+\lambda_{i}}{1-\lambda_{i}}.\qed$
Lemma 39.
Let $\mathbb{S}=\{s_{1},\dots,s_{n}\}$, $f:\mathbb{S}\to\mathbb{R}$ and $\lambda>1$. It holds for Algorithm 2 with $\mathrm{supp}(\pi)=\mathbb{S}$ that
$\mathrm{var}(P_{\lambda},\bar{f})=\sum_{i\in\{1,\dots,n\}\setminus m}\langle u_{i}|\bar{f}\rangle_{\pi}^{2}\frac{1+\lambda_{i}}{1-\lambda_{i}},$
where $\lambda_{1},\dots,\lambda_{n}$ are eigenvalues and $u_{1},\dots,u_{n}$ eigenvectors of matrix $P_{\lambda}$ with $\lambda_{m}=1$ for some $m\in\{1,\dots,n\}$, $|\lambda_{i}|<1$ for all $i\neq m$, $u_{m}=(1,\dots,1)$ and the vectors $u_{1},\dots,u_{n}$ form the orthonormal eigenbasis of $P_{\lambda}$ in the space $(\mathbb{R}^{n},\langle\cdot|\cdot\rangle_{\pi})$.
Proof.
In the finite discrete case Assumption (A1) holds. Algorithm 2 forms $\pi$-reversible Markov chain by Lemma 26, and $P_{\lambda}(x,\{y\})>0$ for all $x,y\in\mathbb{S}$ by Lemma 49 in Appendix E, hence $P_{\lambda}$ is irreducible and aperiodic. Therefore, the claim follows from Lemma 38. ∎
Appendix C Results about real sequences and their sums and differences
The following lemma is well known and is easily verified by Cauchy-Schwarz:
Lemma 40 (Sedrakyan’s inequality).
Let $n\in\mathbb{N}_{+}$, $a_{1},\dots,a_{n}\in\mathbb{R}$ and $b_{1},\dots,b_{n}>0$, then
$\sum_{k=1}^{n}\frac{a_{k}^{2}}{b_{k}}\geq\frac{(\sum_{k=1}^{n}a_{k})^{2}}{\sum_{l=1}^{n}b_{l}},$
where the equality holds if and only if $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{n})$ are linearly dependent.
The following two lemmas use the following basic property: If $n\in\mathbb{N}_{+}$ and $c_{1},\dots,c_{n+1}\in\mathbb{R}$, then
(17) $\sum_{k=1}^{n+1}\sum_{j=1,j\neq k}^{n+1}c_{j}=\sum_{k=1}^{n+1}\sum_{j=1}^{n+1}c_{j}\mathbf{1}\left(j\neq k\right)=\sum_{j=1}^{n+1}c_{j}\sum_{k=1}^{n+1}\mathbf{1}\left(j\neq k\right)=n\sum_{j=1}^{n+1}c_{j}.$
Lemma 41.
Let $n\in\mathbb{N}_{+}$, $a_{1},\dots,a_{n+1}\in\mathbb{R}$, $b_{1},\dots,b_{n+1}\geq 0$, $\sum_{j=1}^{n+1}b_{j}>0$, then
$\frac{1}{n}\sum_{k\in J}\frac{\big(\sum_{i=1,i\neq k}^{n+1}b_{i}a_{i}\big)^{2}}{\sum_{j=1,j\neq k}^{n+1}b_{j}}\geq\frac{\big(\sum_{i=1}^{n+1}b_{i}a_{i}\big)^{2}}{\sum_{j=1}^{n+1}b_{j}},$
where $J=\big\{k\in\{1,2,\dots,{n+1}\}:\sum_{i=1,i\neq k}^{n+1}b_{i}>0\big\}$. Equality holds if and only if $A=(A_{1},\dots,A_{n+1})$ and $B=(B_{1},\dots,B_{n+1})$ are linearly dependent, where $A_{k}=\sum_{j=1,j\neq k}^{n+1}b_{j}a_{j}$ and $B_{k}=\sum_{j=1,j\neq k}^{n+1}b_{j}$.
Proof.
By Sedrakyan’s inequality (Lemma 40)
$\frac{1}{n}\sum_{k\in J}\frac{A_{k}^{2}}{B_{k}}$ $\geq\frac{1}{n}\frac{\big(\sum_{k\in J}A_{k}\big)^{2}}{\sum_{l\in J}B_{l}}=\frac{1}{n}\frac{\big(\sum_{k=1}^{n+1}A_{k}\big)^{2}}{\sum_{l=1}^{n+1}B_{l}},$
for which equality holds if and only if $(A_{k})_{k\in J}$ and $(B_{k})_{k\in J}$ are linearly dependent which is equivalent to $A$ and $B$ being linearly dependent, since $A_{k}=0=B_{k}$ for $k\notin J$. Therefore, the claim follows from (17). ∎
Lemma 42.
Let $n\in\mathbb{N}_{+}$, $a_{1},\dots,a_{n+1}\in\mathbb{R}$, $b_{1},\dots,b_{n+1}\geq 0$, $A_{k}=\sum_{j=1,j\neq k}^{n+1}b_{j}a_{j}$ and $B_{k}=\sum_{j=1,j\neq k}^{n+1}b_{j}$. Then $A=(A_{1},\dots,A_{n+1})$ and $B=(B_{1},\dots,B_{n+1})$ are linearly independent if $a_{i}\neq a_{j}$ and $b_{i},b_{j}>0$ for some $i,j\in\{1,2,\dots,n+1\}$.
Proof.
Assume $A$ and $B$ are linearly dependent i.e. $cA+B=0$ for some $c\neq 0$, then by (17)
$0=\sum_{k=1}^{n+1}(cA_{k}+B_{k})=n\sum_{l=1}^{n+1}b_{l}(ca_{l}+1)$
and thus for all $k\in\{1,2,\dots,n+1\}$
$b_{k}(ca_{k}+1)=\sum_{l=1}^{n+1}b_{l}(ca_{l}+1)-\sum_{m=1,m\neq k}^{n+1}b_{m}(ca_{m}+1)=-(cA_{k}+B_{k})=0,$
hence for each $k$ has to hold either $b_{k}=0$ or $ca_{k}+1=0$. Since $a_{i}\neq a_{j}$ and $b_{i},b_{j}>0$, we get a contradiction. ∎
The final result is about differences of partial sums of sequences.
Lemma 43.
Let $N\in\mathbb{N}_{+}$, $a_{1},a_{2},\dots\in\mathbb{R}$, $b_{1},b_{2},\dots\in\mathbb{R}$, $A_{n}=\sum_{i=1}^{n}a_{i}$ and $B_{n}=\sum_{i=1}^{n}b_{i}$, then
$B_{N+1}B_{N+2}-2B_{N}B_{N+2}+B_{N}B_{N+1}=(b_{N+1}-b_{N+2})B_{N+1}+2b_{N+1}b_{N+2}$
and
$B_{N+1}B_{N+2}A_{N}-2B_{N}B_{N+2}A_{N+1}+B_{N}B_{N+1}A_{N+2}$
$=2b_{N+2}(B_{N+1}A_{N}-B_{N}A_{N+1})-B_{N+1}B_{N}(a_{N+1}-a_{N+2})+B_{N+1}A_{N}(b_{N+1}-b_{N+2}).$
Proof.
These are straightforward, since $B_{n+1}=B_{n}+b_{n+1}$ and $A_{n+1}=A_{n}+a_{n+1}$:
$B_{N+1}B_{N+2}-2B_{N}B_{N+2}+B_{N}B_{N+1}$ $=B_{N+2}b_{N+1}-B_{N}b_{N+2}$
$=B_{N+1}(b_{N+1}-b_{N+2})+2b_{N+1}b_{N+2}$
and thus
$B_{N+1}B_{N+2}A_{N}-2B_{N}B_{N+2}A_{N+1}+B_{N}B_{N+1}A_{N+2}$
$=B_{N+1}A_{N}(b_{N+1}-b_{N+2})+2b_{N+1}b_{N+2}A_{N}$
$\quad-2B_{N}B_{N+2}a_{N+1}+B_{N}B_{N+1}(a_{N+1}+a_{N+2})$
$=B_{N+1}A_{N}(b_{N+1}-b_{N+2})+2(B_{N+1}-B_{N})b_{N+2}A_{N}$
$\quad-2B_{N}(B_{N+1}+b_{N+2})a_{N+1}+B_{N}B_{N+1}a_{N+1}+B_{N}B_{N+1}a_{N+2}$
$=B_{N+1}A_{N}(b_{N+1}-b_{N+2})+2B_{N+1}b_{N+2}A_{N}-2B_{N}b_{N+2}A_{N+1}$
$\quad-B_{N}B_{N+1}(a_{N+1}-a_{N+2}).\qed$
Appendix D Details about modification
In the following lemma the assumption in case (i) is generalisation of sequential convexity and matches this if $a_{i}=i$ for all $i$.
Lemma 44.
Let $(a_{i})_{i\in\mathbb{Z}}\in\mathbb{R}$ be strictly increasing and let us denote $d_{i}=a_{i+1}-a_{i}$. Let $(b_{i})_{i\in\mathbb{Z}}\in\mathbb{R}$ and $f:\mathbb{R}\to\mathbb{R}$,
$f(x)=\Big(1-\frac{x-a_{g(x)}}{a_{g(x)+1}-a_{g(x)}}\Big)b_{g(x)}+\frac{x-a_{g(x)}}{a_{g(x)+1}-a_{g(x)}}b_{g(x)+1},$
where
$g(x)=\inf\{i\in\mathbb{R}:x\geq a_{i}\}.$
The function $f$ is continuous and for all $i\in\mathbb{Z}$ holds $f(a_{i})=b_{i}$. Additionally, if for all $n\in\mathbb{Z}$ holds
  1. (i)
    $\frac{d_{n+1}}{d_{n}+d_{n+1}}b_{n}-b_{n+1}+\frac{d_{n}}{d_{n}+d_{n+1}}b_{n+2}\geq 0$ $(\leq)$, then $f$ is convex (concave).
  2. (ii)
    $b_{n}-b_{n+1}\geq 0$ $(>)$, then $f$ is decreasing (strictly).
  3. (iii)
    $b_{n}-b_{n+1}\leq 0$ $(<)$, then $f$ is increasing (strictly).
  4. (iv)
    $b_{n}\geq 0$ $(b_{n}>0)$, then $f$ is non-negative (positive).
  5. (v)
    $b_{n}\leq 0$ $(b_{n}<0)$, then $f$ is non-positive (negative).
Proof.
Clearly $f$ is continuous on the intervals $a_{i},a_{i+1})$, $f(a_{i})=b_{i}$ and $\lim_{x\uparrow a_{i+1}}f(x)=b_{i+1}$ for all $i\in\mathbb{Z}$, hence $f:\mathbb{R}\to\mathbb{R}$ is continuous.
Let us assume (i) (convex version) holds, then
$b_{i}$ $\geq\frac{d_{i-2}+d_{i-1}}{d_{i-2}}b_{i-1}-\frac{d_{i-1}}{d_{i-2}}b_{i-2}$
$\geq\frac{d_{i-3}+d_{i-2}+d_{i-1}}{d_{i-3}}b_{i-2}-\frac{d_{i-2}+d_{i-1}}{d_{i-3}}b_{i-3}.$
Thus, by induction for any $m\in\mathbb{N}$, $m\geq 2$ it holds that
(18) $\begin{split}b_{i}&\geq\frac{d_{i-m}+\dots+d_{i-1}}{d_{i-m}}b_{i+1-m}-\frac{d_{i+1-m}+\dots+d_{i-1}}{d_{i-m}}b_{i-m}\\ &=\frac{(a_{i}-a_{i-m})b_{i+1-m}-(a_{i}-a_{i+1-m})b_{i-m}}{a_{i+1-m}-a_{i-m}}\end{split}$
and in similar fashion we have
(19) $\begin{split}b_{i}&\geq\frac{(a_{i+m}-a_{i})b_{i+m-1}-(a_{i+m-1}-a_{i})b_{i+m}}{a_{i+m}-a_{i+m-1}}.\end{split}$
Let us denote $\langle x|y\rangle=x_{1}y_{1}+x_{2}y_{2}$ for $x,y\in\mathbb{R}^{2}$ and let $w^{i}=(a_{i},b_{i})\in\mathbb{R}^{2}$. For every $x\in[a_{i},a_{i+1}]$ we have that $(x,f(x))\in J(w^{i},w^{i+1})$, where we denote by $J$ the line segment. Let $\in\mathbb{R}, y<zy<z$ and $t\in(0,1)$, then
$c_{t}:=(1-t)y+tz=(1-s)a_{k}+sa_{k+1}$
for some $s\in[0,1)$ and $k\in\mathbb{Z}$. The vector $v=(b_{k+1}-b_{k},a_{k}-a_{k+1})$ is perpendicular to the line segment $J(w^{k},w^{k+1})\ni(c_{t},f(c_{t}))$. Let us define convex half space that contains this line segment on its boundary
$A=\{u\in\mathbb{R}^{2}:\langle u|v\rangle\leq\langle w^{k}|v\rangle\}.$
Clearly $\langle w^{k}|v\rangle=a_{k}b_{k+1}-a_{k+1}b_{k}=(w^{k+1}|v)$ and by (18) we have for all $i>k+1$ that
$\langle w^{i}|v\rangle\leq a_{i}(b_{k+1}-b_{k})+\frac{(a_{i}-a_{k})b_{k+1}-(a_{i}-a_{k+1})b_{k}}{a_{k+1}-a_{k}}(a_{k}-a_{k+1})=(w^{k}|v).$
Similarly, by (19) we have for all $i<k$ that $\langle w^{i}|v\rangle\leq\langle w^{k}|v\rangle$. Together these imply $w^{i}\in A$ for every $i\in\mathbb{Z}$. If $x<f(c_{t})$, then $(c_{t},x)\notin A$, since
$\langle(c_{t},x)\big|v\rangle$ $=(b_{k+1}-b_{k})c_{t}+(a_{k}-a_{k+1})x>(b_{k+1}-b_{k})c_{t}+(a_{k}-a_{k+1})f(c_{t})$
$=a_{k}b_{k+1}-a_{k+1}b_{k}=\langle w^{k}|v\rangle.$
Since $A$ is convex, we have $(y,f(y)),(z,f(z))\in A$ and $((1-t)y+tz,(1-t)f(y)+tf(z))\in A$, hence $(1-t)f(y)+tf(z)\geq f((1-t)y+tz)$. Thus, we obtain convexity. Concavity follows similarly.
Assume $b_{n}\geq b_{n+1}$ and $x<y$. If $a_{k}\leq x<y<a_{k+1}$ for some $k$, then
$f(x)-f(y)=\frac{y-x}{a_{k+1}-a_{k}}b_{k}+\frac{x-y}{a_{k+1}-a_{k}}b_{k+1}\geq 0$
and $f(x)\geq f(a_{k+1})$ is obvious, thus the claim follows. Similarly, we obtain strict inequality if $b_{i}>b_{i+1}$, which proves (ii). We obtain (iii) similarly. Both (iv) and (v) are clear by the definition of $f$. ∎
Lemma 45.
Let $1\leq a<\infty$. If $g:\mathbb{N}_{+}\to\mathbb{R}$ is a function such that for all $\mathbb{N}\in\mathbb{N}_{+}$
$0\leq g(N)\leq\frac{2a-1}{2a+N-2},$
then for $h:[1,\infty)\to\mathbb{R}$,
$h(\lambda)=(\lfloor\lambda\rfloor+1-\lambda)g(\lfloor\lambda\rfloor)+(\lambda-\lfloor\lambda\rfloor)g(\lfloor\lambda\rfloor+1)$
it holds for all $\lambda\geq 1$ that
$0\leq h(\lambda)\leq\frac{2a}{2a+\lambda-1}.$
Proof.
If $\lambda\geq 1$, $\lambda\in[N,N+1)$, where $N\in\mathbb{N}_{+}$
$h(\lambda)$ $=(N+1-\lambda)g(N)+(\lambda-N)g(N+1)$
$\leq(N+1-\lambda)\frac{2a-1}{2a+N-2}+(\lambda-N)\frac{2a-1}{2a+N-1}$
$=\frac{2a-1}{2a+N-2}\frac{2a+2N-\lambda-1}{2a+N-1}\leq\frac{2a}{2a+N-1}\frac{2a+N-1}{2a+\lambda-1}=\frac{2a}{2a+\lambda-1}.\qed$
Lemma 46.
Let b>1b>-1 and $f:\mathbb{N}_{+}\to\mathbb{R}$, $f(N)=(b+N)^{-1}$, then for $k:[1,\infty)\to\mathbb{R},$
$k(\lambda)=(\lfloor\lambda\rfloor+1-\lambda)f(\lfloor\lambda\rfloor)+(\lambda-\lfloor\lambda\rfloor)f(\lfloor\lambda\rfloor+1)$
it holds for all $\lambda\geq 1$ that
$k(\lambda)$ $=\frac{b+2\lfloor\lambda\rfloor-\lambda+1}{(b+\lfloor\lambda\rfloor+1)(b+\lfloor\lambda\rfloor)}.$
Proof.
We have
$k(\lambda)$ $=(\lfloor\lambda\rfloor+1-\lambda)\frac{1}{b+\lfloor\lambda\rfloor}+(\lambda-\lfloor\lambda\rfloor)\frac{1}{b+\lfloor\lambda\rfloor+1}$
$=(\lfloor\lambda\rfloor+1-\lambda)\frac{b+\lfloor\lambda\rfloor+1}{(b+\lfloor\lambda\rfloor+1)(b+\lfloor\lambda\rfloor)}+(\lambda-\lfloor\lambda\rfloor)\frac{b+\lfloor\lambda\rfloor}{(b+\lfloor\lambda\rfloor+1)(b+\lfloor\lambda\rfloor)}$
$=\frac{b+\lfloor\lambda\rfloor+1}{(b+\lfloor\lambda\rfloor+1)(b+\lfloor\lambda\rfloor)}+\frac{\lfloor\lambda\rfloor-\lambda}{(b+\lfloor\lambda\rfloor+1)(b+\lfloor\lambda\rfloor)}=\frac{b+2\lfloor\lambda\rfloor-\lambda+1}{(b+\lfloor\lambda\rfloor+1)(b+\lfloor\lambda\rfloor)}.\qed$
Appendix E Auxiliaries for the ergodicity of i-SIR
Here we list some properties related to the transition probability of Algorithm 2. The total variation distance between two probabilities $\mu$ and $\nu$ is denoted by
$d_{\mathrm{tv}}(\mu,\nu)=\frac{1}{2}\sup_{\|f\|_{\infty}\leq 1}|\mu(f)-\nu(f)|=\sup_{A}|\mu(A)-\nu(A)|.$
Lemma 47.
Let $\in\mathbb{N}_{+}, M<NM<N$, then
$|P_{M}(x,A)-P_{N}(x,A)|\leq\frac{N-M}{N-1}.$
Proof.
Let us denote $z_{1}=x$, $B_{n}=\sum_{i=1}^{n}w(z_{i})\mathbf{1}\left(z_{i}\in A\right)$ and $C_{n}=\sum_{i=1}^{n}w(z_{i})$, then
$|P_{M}(x,A)-P_{N}(x,A)|$
$\quad=\bigg|\int_{\mathbb{W}_{M}\times\mathbb{X}^{N-M}}\frac{B_{M}C_{N}-B_{N}C_{M}}{C_{M}C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\bigg|$
$\quad=\bigg|\int_{\mathbb{W}_{M}\times\mathbb{X}^{N-M}}\frac{\sum_{i=1}^{M}\sum_{j=M+1}^{N}w(z_{i})w(z_{j})[\mathbf{1}\left(z_{i}\in A\right)-\mathbf{1}\left(z_{j}\in A\right)]}{C_{M}C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\bigg|$
$\quad\leq\int_{\mathbb{W}_{M}\times\mathbb{X}^{N-M}}\frac{\sum_{i=1}^{M}\sum_{j=M+1}^{N}w(z_{i})w(z_{j})}{C_{M}C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
$\quad=\int_{\mathbb{W}_{M}\times\mathbb{X}^{N-M}}\frac{\sum_{j=M+1}^{N}w(z_{j})}{C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)=\int_{\mathbb{W}_{N}}\frac{\sum_{j=M+1}^{N}w(z_{j})}{C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
$\quad=\frac{N-M}{N-1}\int_{\mathbb{W}_{N}}\frac{\sum_{i=2}^{N}w(z_{i})}{C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\leq\frac{N-M}{N-1},$
where the last equality follows since $z_{2},\dots,z_{N}$ are exchangeable under the integral. ∎
Lemma 48.
It holds for all $\lambda,\lambda^{\prime}\in[1,\infty)$ and $x\in\mathbb{S}$ that
$d_{\mathrm{tv}}\big(P_{\lambda}(x,\;\cdot\;),P_{\lambda^{\prime}}(x,\;\cdot\;)\big)\leq\frac{6|\lambda-\lambda^{\prime}|}{\max\{\lambda,\lambda^{\prime}\}}.$
Proof.
Let $N,N^{\prime}\in\mathbb{N}_{+}$. If $N\leq\lambda<\lambda^{\prime}\leq N+1$, then $1/N\leq 2/\lambda^{\prime}$ and
$d_{\mathrm{tv}}\big(P_{\lambda^{\prime}}(x,\;\cdot\;),P_{\lambda}(x,\;\cdot\;)\big)$
$\quad=\sup_{A}|P_{\lambda^{\prime}}(x,A)-P_{\lambda}(x,A)|=\sup_{A}|(\lambda-\lambda^{\prime})P_{N}(x,A)+(\lambda^{\prime}-\lambda)P_{N+1}(x,A)|$
$\quad=|\lambda^{\prime}-\lambda|\sup_{A}|P_{N+1}(x,A)-P_{N}(x,A)|\leq\frac{1}{N}|\lambda^{\prime}-\lambda|,$
where the last inequality follows from Lemma 47. If $N<\lambda\leq N+1\leq N^{\prime}<\lambda^{\prime}\leq N^{\prime}+1$, then by the previous inequality, Lemma 47 and triangle inequality
$d_{\mathrm{tv}}\big(P_{N^{\prime}}(x,\;\cdot\;),P_{N}(x,\;\cdot\;)\big)$
$\quad\leq d_{\mathrm{tv}}\big(P_{\lambda^{\prime}}(x,\;\cdot\;),P_{N^{\prime}}(x,\;\cdot\;)\big)+d_{\mathrm{tv}}\big(P_{N^{\prime}}(x,\;\cdot\;),P_{N+1}(x,\;\cdot\;)\big)+d_{\mathrm{tv}}\big(P_{N+1}(x,\;\cdot\;),P_{\lambda}(x,\;\cdot\;)\big)$
$\quad\leq|\lambda^{\prime}-N^{\prime}|\frac{1}{N^{\prime}}+|N^{\prime}-(N+1)|\frac{1}{N^{\prime}-1}+|N+1-\lambda|\frac{1}{N}\leq\frac{\lambda^{\prime}-(N+1)}{N^{\prime}-1}+\frac{N+1-\lambda}{N}$
$\quad=\frac{N(\lambda^{\prime}-\lambda)+(N^{\prime}-1-N)(N+1-\lambda)}{N(N^{\prime}-1)}\quad\leq\frac{N(\lambda^{\prime}-\lambda)+N^{\prime}-(N+1)}{N(N^{\prime}-1)}$
$\quad\leq\frac{N(\lambda^{\prime}-\lambda)+\lambda^{\prime}-\lambda}{N(N^{\prime}-1)}=\frac{(N+1)(\lambda^{\prime}-\lambda)}{N(N^{\prime}-1)}\leq\frac{2(\lambda^{\prime}-\lambda)}{N^{\prime}-1}\leq\frac{6(\lambda^{\prime}-\lambda)}{N^{\prime}+1}\leq\frac{6(\lambda^{\prime}-\lambda)}{\lambda^{\prime}}.$
Therefore, we obtain the claim. ∎
Lemma 49.
If Assumption (A1) holds, then for all $x\in\mathbb{S}$ and $\lambda\geq 1$
$\frac{\lambda-1}{2\hat{w}+\lambda-1}\pi(A)$ $\leq P_{\lambda}(x,A)\leq\pi(A)+\pi(A^{c})\frac{2\hat{w}}{2\hat{w}+\lambda-1}.$
Proof.
Let us denote $z_{1}=x$. Let $N\in\mathbb{N}_{+}$, $N\geq 2$, then by dropping the first term $i=1$ from (1) an application of Jensen’s inequality yields [e.g. andrieu-lee-vihola]
$P_{N}(z_{1},A)$ $\geq\frac{N-1}{2\hat{w}+N-2}\pi(A),$
which also holds for $N=1$, since $P_{1}(z_{1},A)\geq 0$. If $\lambda\geq 1$, $N\in\mathbb{N}_{+}$, $N\leq\lambda<N+1$, then
$P_{\lambda}(z_{1},A)$ $\geq(N+1-\lambda)\pi(A)\frac{N-1}{2\hat{w}+N-2}+(\lambda-N)\pi(A)\frac{N}{2\hat{w}+N-1}$
$\geq\pi(A)(N+1-\lambda)\frac{N-1}{2\hat{w}+N-1}+\pi(A)(\lambda-N)\frac{N}{2\hat{w}+N-1}$
$=\pi(A)\frac{\lambda-1}{2\hat{w}+N-1}\geq\pi(A)\frac{\lambda-1}{2\hat{w}+\lambda-1}.$
Let $N\in N_{+}$, $N\geq 2$, then
$P_{N}(z_{1},A)$
$\quad=\int_{\mathbb{X}^{N-1}}\frac{\sum_{i=1}^{N}w(z_{i})\mathbf{1}\left(z_{i}\in A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})=1-\int_{\mathbb{X}^{N-1}}\frac{\sum_{i=1}^{N}w(z_{i})\mathbf{1}\left(z_{i}\notin A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})$
$\quad=1-(N-1)\int_{\mathbb{X}^{N-1}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})$
$\quad\leq 1-(N-1)\int_{\mathbb{X}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{w(z_{1})+w(z_{2})+N-2}q(\mathrm{d}z_{2})\leq 1-(N-1)\int_{\mathbb{X}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{2\hat{w}+N-2}q(\mathrm{d}z_{2})$
$\quad=1-(N-1)\frac{\pi(A^{c})}{2\hat{w}+N-2}=\pi(A)+\pi(A^{c})\frac{2\hat{w}-1}{2\hat{w}+N-2},$
where the first inequality follows from Jensen’s inequality. The case $N=1$ is clear. It follows by Lemma 45 in Appendix D, that for all $\lambda\geq 1$
$0\leq P_{\lambda}(z_{1},\{A\})\leq\pi(A)+\pi(A^{c})\frac{2\hat{w}}{2\hat{w}+\lambda-1}.\qed$
Lemma 50.
Let $s>1$. If Assumption (A1) holds, then there exist constants $C<\infty$ and $\rho\in[0,1)$ such that for all $x\in\mathbb{S}$, $\lambda\in[s,\infty)$ and $k\geq 0$ holds
$d_{\mathrm{tv}}(P_{\lambda}^{k}(x,\;\cdot\;),\pi)\leq C\rho^{k}.$
Proof.
By the lower bound in Lemma 49 we have for all $x\in\mathbb{S}$ and $\lambda\geq s$ a constant $c=\frac{s-1}{2\hat{w}+s-1}\in(0,1)$ such that
$P_{\lambda}(x,A)=c\pi(A)+(1-c)Q_{\lambda}(x,A),$
where $Q_{\lambda}$ is probability kernel. For probability kernel $R$ and signed measure $\nu$ on a measurable space $(\mathbb{X},\mathcal{X})$ we have by Hahn decomposition theorem $\nu$-positive set $B^{+}$ and $\nu$-negative set $B^{-}$ for which $\mathbb{X}=B^{+}\cup B^{-}$ and if $G\in\mathcal{X}$, then
$|\nu R(G)|$ $=|\int\nu(\mathrm{d}x)R(x,G)|=|\int_{B^{+}}\nu(\mathrm{d}x)R(x,G)+\int_{B^{-}}\nu(\mathrm{d}x)R(x,G)|$
$\leq\max\Big\{\Big|\int_{B^{+}}\nu(\mathrm{d}x)R(x,G)\Big|,\Big|\int_{B^{-}}\nu(\mathrm{d}x)R(x,G)\Big|\Big\}\leq\max\{|\nu(B^{+})|,|\nu(B^{-})|\}$
$=\sup_{A}|\nu(A)|$
and thus
$\sup_{A}|\nu R(A)|\leq\sup_{A}|\nu(A)|,$
which ensures for $x\in\mathbb{S}$ and $k\geq 1$
$\sup_{A}|P_{\lambda}^{k}(x,A)-\pi(A)|$
$\quad=\sup_{A}|\delta_{x}P_{\lambda}^{k}(A)-\pi(A)|=\sup_{A}|\delta_{x}P_{\lambda}^{k-1}P_{\lambda}(A)-\pi P_{\lambda}^{k-1}P_{\lambda}(A)|$
$\quad=\sup_{A}|c\pi(A)+(1-c)\delta_{x}P_{\lambda}^{k-1}Q_{\lambda}(A)-c\pi(A)-(1-c)\pi P_{\lambda}^{k-1}Q_{\lambda}(A)|$
$\quad=(1-c)\sup_{A}|\delta_{x}P_{\lambda}^{k-1}Q_{\lambda}(A)-\pi P_{\lambda}^{k-1}Q_{\lambda}(A)|\leq(1-c)\sup_{A}|\delta_{x}P_{\lambda}^{k-1}(A)-\pi P_{\lambda}^{k-1}(A)|$
$\quad\leq\dotsi$
$\quad\leq(1-c)^{k}\sup_{A}|\delta_{x}Q_{\lambda}(A)-\pi Q_{\lambda}(A)|\leq(1-c)^{k}\sup_{A}|\delta_{x}(A)-\pi(A)|\leq(1-c)^{k}.\qed$
Appendix F Cost function
The following holds for affine cost function:
Proposition 51.
Let $c,u:(1,\infty)\to\mathbb{R}$, $c(\lambda)=a+b\lambda$ and $u(\lambda)=c(\lambda)\frac{d+\lambda-1}{\lambda-1}$, where $a\geq 0$, $b>0$ and $d>0$, then
  1. (i)
    $u$ is strictly convex positive function,
  2. (ii)
    $\lim_{\lambda\to\infty}u(\lambda)=\lim_{\lambda\to\infty}c(\lambda)=\infty$ and $\lim_{\lambda\to 1_{+}}u(\lambda)=\infty$,
  3. (iii)
    $\lim_{\lambda\to\infty}[u(\lambda)-c(\lambda)]=bd$,
  4. (iv)
    the minimum value of $u$ is $2\sqrt{bd(a+b)}+a+b+bd$ attained at $\sqrt{d(a/b+1)}+1$ and the function $c$ attains this value at $2\sqrt{d(a/b+1)}+d+1$.
Proof.
Let $g_{1},g_{2}:(1,\infty)\to\mathbb{R}$, $g_{1}(\lambda)=\frac{d+\lambda-1}{\lambda-1}$ and $g_{2}(\lambda)=\lambda\frac{d+\lambda-1}{\lambda-1}$, then
$\frac{\partial}{\partial\lambda}g_{1}(\lambda)$ $=-\frac{d}{(\lambda-1)^{2}},\qquad\frac{\partial^{2}}{\partial\lambda^{2}}g_{1}(\lambda)=\frac{2d}{(\lambda-1)^{3}},$
$\frac{\partial}{\partial\lambda}g_{2}(\lambda)$ $=1-\frac{d}{(\lambda-1)^{2}},\qquad\frac{\partial^{2}}{\partial\lambda^{2}}g_{2}(\lambda)=\frac{2d}{(\lambda-1)^{3}}.$
Therefore
$u(\lambda)=ag_{1}(\lambda)+bg_{2}(\lambda)>0,\qquad\frac{\partial}{\partial\lambda}u(\lambda)=b-\frac{d(b+a)}{(\lambda-1)^{2}},\qquad\frac{\partial^{2}}{\partial\lambda^{2}}u(\lambda)=\frac{2d(b+a)}{(\lambda-1)^{3}}>0,$
from which we obtain (i). The minimum value of $u$ can be solved from $\frac{\partial}{\partial\lambda}u(\lambda)$, which is at $\lambda_{m}=1+\sqrt{\frac{d(b+a)}{b}}$, thus the minimum value
$m=u(\lambda_{m})=(a+b)\bigg(\sqrt{\frac{bd}{d(a+b)}}+1\bigg)+b\bigg(d+\sqrt{\frac{d(b+a)}{b}}\bigg)=2\sqrt{bd(a+b)}+a+b+bd$
and if $c(\lambda)=m$, then $\lambda=\frac{m-a}{b}$ and we have (iv). We obtain (iii) from
$u(\lambda)-c(\lambda)=\frac{d(a+b\lambda)}{\lambda-1}\xrightarrow{\lambda\to\infty}bd,$
and (ii) is clear by the definitions. ∎
Lemma 52.
Let Assumption (A1) hold, $f\in L_{0,+}^{2}(\pi)$ and $\lambda_{1},\lambda_{2}\in(1,2]$, $\lambda_{1}<\lambda_{2}$, then
$\lambda_{1}\cdot\mathrm{var}(P_{\lambda_{1}},f)-\lambda_{2}\cdot\mathrm{var}(P_{\lambda_{2}},f)>0.$
Proof.
Let $g\in L_{0}^{2}(\pi)$, then
$\mathcal{E}_{\lambda_{1}}(g)$ $=(\lambda_{1}-\lambda_{2})\mathcal{E}_{2}(g)+\mathcal{E}_{\lambda_{2}}(g).$
By Lemma 26, we can set $h_{\lambda_{2}}=(I-P_{\lambda_{2}})^{-1}f$, for which
$\mathrm{var}(P_{\lambda_{2}},f)=2\langle f|h_{\lambda_{2}}\rangle_{\pi}-\langle f|f\rangle_{\pi},\qquad\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})=\langle h_{\lambda_{2}}|f\rangle_{\pi}=\langle f|h_{\lambda_{2}}\rangle_{\pi},$
and by Lemma 60 in Appendix H
$\mathrm{var}(P_{\lambda_{1}},f)=2\sup_{g\in L_{0}^{2}(\pi)}[2\langle f|g\rangle_{\pi}-\mathcal{E}_{\lambda_{1}}(g)]-\langle f|f\rangle_{\pi}\geq 2[2\langle f|h_{\lambda_{2}}\rangle_{\pi}-\mathcal{E}_{\lambda_{1}}(h_{\lambda_{2}})]-\langle f|f\rangle_{\pi}.$
Therefore
$\lambda_{1}\cdot\mathrm{var}(P_{\lambda_{1}},f)-\lambda_{2}\cdot\mathrm{var}(P_{\lambda_{2}},f)$
$\quad\geq\lambda_{1}2[2\langle f|h_{\lambda_{2}}\rangle_{\pi}-\mathcal{E}_{\lambda_{1}}(h_{\lambda_{2}})]-\lambda_{2}2\langle f|h_{\lambda_{2}}\rangle_{\pi}+(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}$
$\quad=\lambda_{1}2[2\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})-(\lambda_{1}-\lambda_{2})\mathcal{E}_{2}(h_{\lambda_{2}})-\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})]-2\lambda_{2}\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})+(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}$
$\quad=\lambda_{1}2(\lambda_{2}-\lambda_{1})\mathcal{E}_{2}(h_{\lambda_{2}})-2(\lambda_{2}-\lambda_{1})\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})+(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}$
$\quad\geq 2(\lambda_{2}-\lambda_{1})[\mathcal{E}_{2}(h_{\lambda_{2}})-\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})]+(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}$
$\quad\geq(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}>0,$
where the second last inequality follows from Theorem 20 and the fact that $\lambda_{1}>1$. ∎
For Algorithm 2, derivatives of holding probability and $\psi$ (given in (14)) are defined for non-integer $\lambda>1$,
$\frac{\partial}{\partial\lambda}P_{\lambda}(x,\{x\})$ $=P_{\lfloor\lambda+1\rfloor}(x,\{x\})-P_{\lfloor\lambda\rfloor}(x,\{x\}),\qquad\psi^{\prime}(\lambda)=\psi(\lfloor\lambda+1\rfloor)-\psi(\lfloor\lambda\rfloor).$
Lemma 53.
Let $g,h:(1,\infty)\to\mathbb{R}$,
$g(\lambda)=\lambda\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)},\qquad h(\lambda)=\lambda\frac{1+\psi(\lambda)}{1-\psi(\lambda)},$
then $g$ is strictly decreasing and strictly convex when $\lambda\in(1,2]$, same holds for $h$ if $(\mathbb{S})>1$.
Proof.
We have for $\lambda>1$, that $0<\epsilon(\lambda)<1$, and $\epsilon$ is decreasing by Theorem 21. Same properties hold for $\psi$ when $(\mathbb{S})>1$ by Lemma 25. We have for $\lambda\in(1,2)$, that
$g^{\prime}(\lambda)=\frac{2\lambda\epsilon^{\prime}(\lambda)+1-\epsilon(\lambda)^{2}}{(1-\epsilon(\lambda))^{2}},$
where
$2\lambda\epsilon^{\prime}(\lambda)+1-\epsilon(\lambda)^{2}$ $=2\lambda\big(\epsilon(2)-1\big)+1-\big((2-\lambda)+(\lambda-1)\epsilon(2)\big)^{2}$
$=2\lambda\big(\epsilon(2)-1\big)+2(\lambda-1)\big(1-\epsilon(2)\big)-(\lambda-1)^{2}\big(1-\epsilon(2)\big)^{2}$
$=-2\big(1-\epsilon(2)\big)-(\lambda-1)^{2}\big(1-\epsilon(2)\big)^{2}<0,$
and
$g^{\prime\prime}(\lambda)$ $=\frac{(1-\epsilon(\lambda))[2\epsilon^{\prime}(\lambda)-2\epsilon(\lambda)\epsilon^{\prime}(\lambda)]+2\epsilon^{\prime}(\lambda)[2\lambda\epsilon^{\prime}(\lambda)+1-\epsilon(\lambda)^{2}]}{(1-\epsilon(\lambda))^{3}}$
$=-\frac{2\epsilon^{\prime}(\lambda)\big[-(1-\epsilon(\lambda))^{2}+2\big(1-\epsilon(2)\big)+(\lambda-1)^{2}\big(1-\epsilon(2)\big)^{2}\big]}{(1-\epsilon(\lambda))^{3}}$
$\geq-\frac{2\epsilon^{\prime}(\lambda)\big[\big(1-\epsilon(2)\big)+(\lambda-1)^{2}\big(1-\epsilon(2)\big)^{2}\big]}{(1-\epsilon(\lambda))^{3}}>0.$
Since $g$ is continuous, we get the claim. The proof is similar for $h$. ∎
When we combine the affine cost function with functions in (15), we obtain the following proposition:
Proposition 54.
Let $c:(1,\infty)\to\mathbb{R}$, $c(\lambda)=a+b\lambda$, where $a\geq 0$ and $b>0$. Let Assumption (A1) hold and $f\in L_{0,+}^{2}(\pi)$, then $\lambda\mapsto c(\lambda)V_{f}(\lambda)$, $\lambda\mapsto c(\lambda)G_{f}(\lambda)$ and $\lambda\mapsto c(\lambda)H_{f}(\lambda)$ are strictly decreasing when $\lambda\in(1,2]$.
Proof.
Follows from Theorem 27, Lemma 24 and 25 combined with Lemma 52 and 53. ∎
Appendix G Finite discrete case for i-SIR
Estimating the asymptotic variance of i-SIR can be challenging. In the finite discrete case asymptotic variance can be solved via the formula in Lemma 39 in Appendix B, since the transition probability matrix $P_{\lambda}$ as well as the eigenvalues and eigenvectors can be solved, even tough this can be extremely slow for large $\lambda$ when $\mathbb{S}$ has many states. However, in the finite case transition probability matrix $P_{N}$ can be expressed with the help of multinomial distribution:
Proposition 55.
Let us have Algorithm 1 with finite discrete $\pi$, $\mathbb{S}=\{s_{1},s_{2},\dots,s_{n}\}$, $n\in\mathbb{N}_{+}$ and let us denote $U=\mathrm{supp}(q)\setminus\mathbb{S}$.
  1. (i)
    If $U=\emptyset$, let $Q=(q(\{s_{1}\}),\dots,q(\{s_{n}\}))$.
  2. (ii)
    If $U\neq\emptyset$, let $Q=(q(\{s_{1}\}),\dots,q(\{s_{n}\}),q(U))$.
For both cases (i) and (ii) we have for $N\in\mathbb{N}$, $N\geq 2$ and $i,j\in\{1,\dots,n\}$
(20) $P_{N}(s_{i},\{s_{j}\})$ $=\mathbb{E}\Big[(\mathbf{1}\left(i=j\right)+Z_{j})\frac{w(s_{j})}{w(s_{i})+\sum_{k=1}^{n}Z_{k}w(s_{k})}\Big],$
(21) $\epsilon(N,s_{i})$ $=\mathbb{E}\Big[\frac{w(s_{i})}{w(s_{i})+\sum_{k=1}^{n}Z_{k}w(s_{k})}\Big],$
$\epsilon(N)$ $=\mathbb{E}\Big[\frac{w(Y)}{w(Y)+\sum_{k=1}^{n}Z_{k}w(s_{k})}\Big],$
where $Z\sim\mathrm{Multinom}(N-1,Q)$ and $Y\sim\pi$ are independent.
Proof.
Let us first take $U=\emptyset$ and $Q=(q(\{s_{1}\}),\dots,q(\{s_{n}\}))$. For Algorithm 1 we have $Y_{1}^{1}=X_{0}$, $Y_{1}^{2:N}\sim q$ and $I\sim\mathrm{Categorical}(W_{1}^{1:N})$, where $Y^{2:N}$ are independent of each other. Therefore
$Z_{1}=\Big(\sum_{j=2}^{N}\mathbf{1}\left(Y_{1}^{j}=s_{1}\right),\dots,\sum_{j=2}^{N}\mathbf{1}\left(Y_{1}^{j}=s_{n}\right)\Big)\sim\mathrm{Multinom}(N-1,q).$
Let $v\in\mathbb{R}^{n}$ and let us define
$A_{1}^{v}=\{Z_{1}=v\},\qquad B=\Big\{v\in\mathbb{N}^{n}:\sum_{k=1}^{n}v_{k}=N-1\Big\}.$
Then for $i,j\in\{1,\dots,n\}$
$\mathbb{P}(X_{1}=s_{j}\mid X_{0}=s_{i})$ $=\sum_{v\in B}\mathbb{P}(X_{1}=s_{j},A_{1}^{v}\mid X_{0}=s_{i})$
$=\sum_{v\in B}\mathbb{P}(X_{1}=s_{j}\mid X_{0}=s_{i},A_{1}^{v})\mathbb{P}(A_{1}^{v})$
$=\sum_{v\in B}\sum_{k=1}^{N}\mathbb{P}(X_{1}=s_{j},I=k\mid X_{0}=s_{i},A_{1}^{v})\mathbb{P}(A_{1}^{v})$
$=\begin{cases}\sum_{v\in B}(1+v_{i})\frac{w(s_{i})}{w(s_{i})+\sum_{k=1}^{n}v_{k}w(s_{k})}\mathbb{P}(A_{1}^{v}),&i=j\\ \sum_{v\in B}v_{j}\frac{w(s_{j})}{w(s_{i})+\sum_{k=1}^{n}v_{k}w(s_{k})}\mathbb{P}(A_{1}^{v}),&i\neq j.\end{cases}$
For the rejection
$\mathbb{P}(I=1\mid X_{0}=s_{i})$ $=\sum_{v\in B}\mathbb{P}(I=1,A_{1}^{v}\mid X_{0}=s_{i})=\sum_{v\in B}\mathbb{P}(I=1\mid X_{0}=s_{i},A_{1}^{v})\mathbb{P}(A_{1}^{v})$
$=\sum_{v\in B}\frac{w(s_{i})}{w(s_{i})+\sum_{k=1}^{n}v_{k}w(s_{k})}\mathbb{P}(A_{1}^{v}).$
Since $(X_{i})$ is Markov chain we obtain the case for $U=\emptyset$. The case $U\neq\emptyset$ follows similarly when we have instead $Q=(q(\{s_{1}\}),\dots,q(\{s_{n}\}),q(U))$, $v\in\mathbb{R}^{n+1}$,
$Z_{1}$ $=\Big(\sum_{j=2}^{N}\mathbf{1}\left(Y_{1}^{j}=s_{1}\right),\dots,\sum_{j=2}^{N}\mathbf{1}\left(Y_{1}^{j}=s_{n}\right),\sum_{j=2}^{N}\mathbf{1}\left(Y_{1}^{j}\notin\mathbb{S}\right)\Big)\sim\mathrm{Multinom}(N-1,Q),$
$A_{1}^{v}$ $=\{Z_{1}=v\},\qquad B=\Big\{v\in\mathbb{N}^{n+1}:\sum_{k=1}^{n+1}v_{k}=N-1\Big\},$
since $w(s)=0$ for all $s\notin\mathbb{S}$. ∎
Proposition 55 can be easily generalised to Algorithm 2 and gives possibilities to estimate the probability transition matrix, holding probability and rejection probabilities with multinomial distribution. We have also the following special case:
Proposition 56.
If $(\mathbb{S})=2$, i.e. $\mathbb{S}=\{s_{1},s_{2}\}$, then it holds for $f\in L_{0,+}^{2}(\pi)$ that
$H_{f}(\lambda)=\mathrm{var}(P_{\lambda},f)=\mathrm{var}_{\pi}(f)\frac{P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})}{2-(P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\}))}.$
Proof.
We have $P_{\lambda}$ is $\pi$-reversible by Lemma 26. We can solve eigenvalues $\lambda_{1}=1$, $\lambda_{2}=P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})-1$ and eigenvectors $v_{1}=(1,1)$, $v_{2}=(-\sqrt{\frac{\pi(\{s_{2}\})}{\pi(\{s_{1}\})}},\sqrt{\frac{\pi(\{s_{1}\})}{\pi(\{s_{2}\})}})$ to match conditions in Lemma 39 in Appendix B and thus
$\mathrm{var}(P_{\lambda},f)$ $=\langle v_{2}|f\rangle_{\pi}^{2}\frac{P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})}{2-(P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\}))}$
$=\mathrm{var}_{\pi}(f)\frac{P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})}{2-[P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})]},$
since
$\langle v_{2}|f\rangle_{\pi}^{2}$ $=\bigg(-\sqrt{\frac{\pi(\{s_{2}\})}{\pi(\{s_{1}\})}}f(s_{1})\pi(\{s_{1}\})+\sqrt{\frac{\pi(\{s_{1}\})}{\pi(\{s_{2}\})}}f(s_{2})\pi(\{s_{2}\})\bigg)^{2}$
$=\pi(\{s_{2}\})\pi(\{s_{1}\})\big(f(s_{1})-f(s_{2})\big)^{2},$
$\mathbb{E}_{\pi}[f]=0$ and
$\mathrm{var}_{\pi}(f)$ $=\pi(\{s_{1}\})[\pi(\{s_{1}\})+\pi(\{s_{2}\})]f(s_{1})^{2}+\pi(\{s_{2}\})[\pi(\{s_{1}\})+\pi(\{s_{2}\})]f(s_{2})^{2}$
$=2\pi(\{s_{1}\})\pi(\{s_{2}\})f(s_{1})f(s_{2})+\pi(\{s_{1}\})\pi(\{s_{2}\})[f(s_{1})-f(s_{2})]^{2}+\sum_{i=1}^{2}\pi(\{s_{i}\})^{2}f(s_{i})^{2}$
$=\pi(\{s_{1}\})\pi(\{s_{2}\})[f(s_{1})-f(s_{2})]^{2}.$
We have $1-\pi(\{s_{1}\})^{2}-\pi(\{s_{2}\})^{2}=2\pi(\{s_{1}\})\pi(\{s_{2}\})$ and by reversibility
$-1+\sum_{i=1}^{2}\pi(\{s_{i}\})P_{\lambda}(s_{i},\{s_{i}\})=-\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{2}\})-\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{1}\})$
$\quad=-[\pi(\{s_{1}\})+\pi(\{s_{2}\})][\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{2}\})+\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{1}\})]$
$\quad=-2\pi(\{s_{1}\})\pi(\{s_{2}\})[P_{\lambda}(s_{1},\{s_{2}\})+P_{\lambda}(s_{2},\{s_{1}\})],$
therefore
$H_{f}(\lambda)$ $=\frac{1-2\pi(\{s_{1}\})^{2}-2\pi(\{s_{2}\})^{2}+\sum_{i=1}^{2}\pi(\{s_{i}\})P_{\lambda}(s_{i},\{s_{i}\})]}{1-\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{1}\})-\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{2}\})}\mathrm{var}_{\pi}(f)$
$=\frac{2-P_{\lambda}(s_{1},\{s_{2}\})-P_{\lambda}(s_{2},\{s_{1}\})]}{P_{\lambda}(s_{1},\{s_{2}\})+P_{\lambda}(s_{2},\{s_{1}\})}\mathrm{var}_{\pi}(f)=\frac{P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})}{2-P_{\lambda}(s_{1},\{s_{1}\})-P_{\lambda}(s_{2},\{s_{2}\})]}\mathrm{var}_{\pi}(f)$
and we obtain the claim. ∎
Appendix H Some properties on Hilbert spaces and reversible Markov chains
On most functional analysis books interesting operator properties are shown for the complex Hilbert space. However as mentioned in [bogachev-smolyanov] many of the results also hold in the real case for self-adjoint operators and can be obtained via complexification. We shall denote by $\|\cdot\|$ the norm induced by the inner product $\langle\cdot|\cdot\rangle$ of a Hilbert space $\mathcal{H}$ that is either real or complex.
The following lemma is combination of Theorems from [rudin, p. 330–331]:
Lemma 57.
Let $A$ be a bounded operator on a complex Hilbert space $\mathcal{H}$ satisfying $\langle f|Af\rangle\geq 0$ for all $f\in\mathcal{H}$. Then it holds that $A^{*}=A$ and there exists square root $A^{1/2}$ for which $\langle f|A^{1/2}f\rangle\geq 0$ for all $f\in\mathcal{H}$. Additionally if $A$ is invertible, so is $A^{1/2}$.
Lemma 58.
Let $A$ be a bounded self-adjoint operator on a Hilbert space $\mathcal{H}$, satisfying $\langle f|Af\rangle\geq 0$ for all $f\in\mathcal{H}$ and such that the inverse $A^{-1}$ exists. Then it also holds that $\langle f|Af\rangle=0$ if and only if $\|f\|=0$.
Proof.
It is clear by Cauchy-Schwarz that $\|f\|=0$ implies $\langle f|Af\rangle=0$. For the reverse it is enough to show that $\|f\|>0$ implies $\langle f|Af\rangle>0$. If $\mathcal{H}$ is complex Hilbert space, then by Theorem 57 there exists invertible square root $A^{1/2}$ for which $(A^{1/2})^{*}=A^{1/2}$ and thus for $f\in\mathcal{H}$, $\|f\|>0$
$\langle f|Af\rangle=\langle f|A^{1/2}A^{1/2}f\rangle=\langle A^{1/2}f|A^{1/2}f\rangle>0,$
where the strict inequality follows since $A^{1/2}$ is invertible. In the case of Real Hilbert space $\mathcal{H}$ let us do complexification of the Hilbert space, then $A_{\mathbb{C}}(g+ih)=Ag+iAh$, $g,h\in\mathcal{H}$. Since $A$ is self-adjoint and $\langle f|Af\rangle\geq 0$ for all $f\in\mathcal{H}$
$\langle g+ih|A_{\mathbb{C}}(g+ih)\rangle_{\mathbb{C}}$ $=\langle g|Ag\rangle-i\langle g|Ah\rangle+i\langle h|Ag\rangle+\langle h|Ah\rangle$
$=\langle g|Ag\rangle-i\langle Ag|h\rangle+i\langle h|Ag\rangle+\langle h|Ah\rangle$
$=\langle g|Ag\rangle+\langle h|Ah\rangle\geq 0.$
Thus, it holds that $\langle f|A_{\mathbb{C}}f\rangle_{\mathbb{C}}\geq 0$ for all $f\in\mathcal{H}_{\mathbb{C}}$ for which the earlier result holds and therefore for all $f\in\mathcal{H}$, $\|f\|>0$
$\langle f|Af\rangle=\langle f+i0|A_{\mathbb{C}}(f+i0)\rangle_{\mathbb{C}}>0.\qed$
Lemma 59.
Let $A$ be a bounded self-adjoint operator on a Hilbert space $\mathcal{H}$, satisfying $\langle f|Af\rangle\geq 0$ for all $f\in\mathcal{H}$, then $A$ is invertible if there exists such $c>0$ that
$\|Af\|\geq c\|f\|$
for all $f\in\mathcal{H}$.
Proof.
Result holds for complex Hilbert space [e.g. rudin, Theorem 12.12]. In the case of Real Hilbert space $\mathcal{H}$ let us do complexification as in proof of Lemma 58. Then $\langle f|A_{\mathbb{C}}f\rangle\geq 0$ for all $f\in\mathcal{H}_{\mathbb{C}}$, and by Lemma 57 $A_{\mathbb{C}}$ is self-adjoint. Additionally for $f,g\in\mathcal{H}$
$\langle f+ig|f+ig\rangle_{\mathbb{C}}$ $=\langle f|f\rangle-i\langle f|g\rangle+i\langle g|f\rangle+\langle g|g\rangle$
$=\langle f|f\rangle-i\langle f|g\rangle+i\langle f|g\rangle+\langle g|g\rangle$
$=\langle f|f\rangle+\langle g|g\rangle,$
hence
$\langle A(f+ig)|A(f+ig)\rangle_{\mathbb{C}}$ $=\langle Af|Af\rangle+\langle Ag|Ag\rangle\geq c^{2}\langle f|f\rangle+c^{2}\langle g|g\rangle=c^{2}\langle f+ig|f+ig\rangle.$
By the complex case result $A_{\mathbb{C}}:\mathcal{H}_{\mathbb{C}}\to\mathcal{H}_{\mathbb{C}}$ is invertible and therefore $A:\mathcal{H}\to\mathcal{H}$ is invertible. ∎
Lemma 60.
Let $A$ be a bounded self-adjoint operator on a Hilbert space $\mathcal{H}$, satisfying $\langle f|Af\rangle\geq 0$ for all $f\in\mathcal{H}$ and such that the inverse $A^{-1}$ exists. Then
$\langle f|A^{-1}f\rangle=\sup_{g\in\mathcal{H}}[2\langle f|g\rangle-\langle g|Ag\rangle],$
where the supremum is attained uniquely with $g=A^{-1}f$ up to $\mathcal{H}$-equivalence class.
Proof.
As in [andrieu-vihola-order, Lemma 16] equalities follow from noticing
(22) $\langle f|A^{-1}f\rangle-[2\langle f|g\rangle-\langle g|Ag\rangle]=\langle A^{-1}f-g,A(A^{-1}f-g)\rangle\geq 0,$
where the equality holds for $g=A^{-1}f$. By Lemma 58, equation (22) is equal to zero if and only if $\|A^{-1}f-g\|=0$. ∎
Assumption (A3).
Let $J\subset\mathbb{R}$ and $(A_{\lambda})_{\lambda\in J}$ be collection of bounded self-adjoint operators on Hilbert space $\mathcal{H}$ for which $\langle f|A_{\lambda}f\rangle\geq 0$ for all $f\in H$, the inverse $A_{\lambda}^{-1}$ exists and $\lambda\mapsto\langle f|A_{\lambda}f\rangle$ is injective for $\|f\|>0$ when $\lambda\in J$.
Lemma 61.
If (A3) holds, $\in J, xyx\neq y$ and $f\in\mathcal{H}$, then $\|A_{x}^{-1}f-A_{y}^{-1}f\|>0$ if $\|f\|>0$.
Proof.
Let us make an assumption that $\|A_{x}^{-1}f-A_{y}^{-1}f\|=0$. Since $A_{\lambda}$ is self-adjoint for all $\lambda\in J$
$\langle A_{y}^{-1}f,A_{y}A_{y}^{-1}f\rangle$ $=\langle f,A_{y}^{-1}f\rangle=\langle A_{x}^{-1}f,A_{x}A_{y}^{-1}f\rangle=\langle A_{y}^{-1}f,A_{x}A_{y}^{-1}f\rangle,$
where the assumption is used on the last equality. We have contradiction, since $\|A_{y}^{-1}f\|>0$ and $\lambda\mapsto\langle A_{y}^{-1}f|A_{x}A_{y}^{-1}f\rangle$ is injective by (A3). ∎
Lemma 62.
If (A3) holds and $\in J, xyx\neq y$, $f\in\mathcal{H}$, $\|f\|>0$, then there is no such $g\in\mathcal{H}$, that both equations
$\sup_{h\in\mathcal{H}}[2\langle f|h\rangle-\langle h|A_{x}h\rangle]=2\langle f|g\rangle-\langle g|A_{x}g\rangle$
and
$\sup_{h\in\mathcal{H}}[2\langle f|h\rangle-\langle h|A_{y}h\rangle]=2\langle f|g\rangle-\langle g|A_{y}g\rangle$
hold.
Proof.
Follows from Lemma 60 and 61. ∎
We have the following result for collection of $\pi$-reversible Markov kernels:
Theorem 63.
Let $J\subset\mathbb{R}$ be and $(Q_{\lambda})_{\lambda\in J}$ be collection of $\pi$-reversible Markov kernels and the inverse operator $(I-Q_{\lambda})^{-1}$ exists for $(I-Q_{\lambda}):L_{0}^{2}(\pi)\to L_{0}^{2}(\pi)$. Let us define for fixed $h\in L_{0}^{2}(\pi)$ functions $u_{h},v_{h}:J\to\mathbb{R}$, $u_{h}(\lambda)=\langle h|Q_{\lambda}h\rangle_{\pi}$, $v_{h}(\lambda)=\mathrm{var}(Q_{\lambda},h)$.
Let us state the following assumptions:
  1. (a)
    $u_{g}$ is decreasing (increasing) function for all $g\in L_{0,+}^{2}(\pi)$.
  2. (b)
    $u_{g}$ is strictly decreasing (increasing) function for all $g\in L_{0,+}^{2}(\pi)$.
  3. (c)
    $u_{g}$ is convex function for all $g\in L_{0,+}^{2}(\pi)$, where $J$ is interval.
  4. (d)
    $u_{g}$ is strictly convex function for all $g\in L_{0,+}^{2}(\pi)$, where $J$ is interval.
For $f\in L_{0,+}^{2}(\pi)$ it holds that
  1. (i)
    $v_{f}$ is decreasing (increasing) function if (a) holds.
  2. (ii)
    $v_{f}$ is strictly decreasing (increasing) function if (b) holds.
  3. (iii)
    $v_{f}$ is convex function if (c) holds.
  4. (iv)
    $v_{f}$ is strictly convex function if (d) holds.
  5. (v)
    $v_{f}$ is strictly convex and strictly decreasing (increasing) function if (b) and (c) hold.
Proof.
Since $Q_{\lambda}$ is $\pi$-reversible, we have $I-Q_{\lambda}:L_{0}^{2}(\pi)\to L_{0}^{2}(\pi)$ is self-adjoint. For $g\in L_{0}^{2}(\pi)$, let us denote the Dirichlet form by
$\mathcal{E}_{\lambda}(g)=\langle g|(I-Q_{\lambda})g\rangle_{\pi}=\langle g|g\rangle_{\pi}-u_{g}(\lambda),$
which is non-negative for the Markov kernel, and the function $\lambda\mapsto-\mathcal{E}_{\lambda}(g)$ preserves conditions in (a)–(d). Let us denote $h_{\lambda}=(I-Q_{\lambda})^{-1}f$. By Lemma 60
$2\langle f|h_{\lambda}\rangle_{\pi}-\mathcal{E}_{\lambda}(h_{\lambda})=\sup_{g\in L_{0}^{2}(\pi)}[2\langle f|g\rangle_{\pi}-\mathcal{E}_{\lambda}(g)]\geq 2\langle f|g\rangle_{\pi}-\mathcal{E}_{\lambda}(g)$
for all $g\in L_{0}^{2}(\pi)$ and
$\mathrm{var}(Q_{\lambda},f)=2\langle f|(I-Q_{\lambda})^{-1}f\rangle_{\pi}-\langle f|f\rangle_{\pi}=2\sup_{g\in L_{0}^{2}(\pi)}[2\langle f|g\rangle_{\pi}-\mathcal{E}_{\lambda}(g)]-\langle f|f\rangle_{\pi}.$
Let $x,y\in J$. When (a) holds with the decreasing case, and $x<y$ (or (a) holds with the increasing case, and $x>y$), then
$\mathrm{var}(Q_{x},f)-\mathrm{var}(Q_{y},f)$ $=2\Big(\sup_{g\in L_{0}^{2}(\pi)}[2\langle f|g\rangle_{\pi}-\mathcal{E}_{x}(g)]-\sup_{h\in L_{0}^{2}(\pi)}[2\langle f|h\rangle_{\pi}-\mathcal{E}_{y}(h)]\Big)$
$\geq 2\Big(\big(2\langle f|h_{y}\rangle_{\pi}-\mathcal{E}_{x}(h_{y})\big)-\big(2\langle f|h_{y}\rangle_{\pi}-\mathcal{E}_{y}(h_{y})\big)\Big)$
$=2\big(-\mathcal{E}_{x}(h_{y})+\mathcal{E}_{y}(h_{y})\big)\geq 0,$
where the last inequality is strict if (b) holds. When (c) holds, $x<y$ and $t\in(0,1)$, then
$\mathrm{var}(Q_{tx+(1-t)y},f)$
$=2\Big(2\langle f|h_{tx+(1-t)y}\rangle_{\pi}-\mathcal{E}_{tx+(1-t)y}(h_{tx+(1-t)y})\Big)-\langle f|f\rangle_{\pi}$
$\leq 2\Big(t[2\langle f|h_{tx+(1-t)y}\rangle_{\pi}-\mathcal{E}_{x}(h_{tx+(1-t)y})]+(1-t)[2\langle f|h_{tx+(1-t)y}\rangle_{\pi}-\mathcal{E}_{y}(h_{tx+(1-t)y})]\Big)-\langle f|f\rangle_{\pi}$
$\leq 2\sup_{g\in L_{0}^{2}(\pi)}\Big(t[2\langle f|g\rangle_{\pi}-\mathcal{E}_{x}(g)]+(1-t)[2\langle f|g\rangle_{\pi}-\mathcal{E}_{y}(g)]\Big)-\langle f|f\rangle_{\pi}$
$\leq 2\Big(t\sup_{g\in L_{0}^{2}(\pi)}[2\langle f|g\rangle_{\pi}-\mathcal{E}_{x}(g)]+(1-t)\sup_{h\in L_{0}^{2}(\pi)}[2\langle f|h\rangle_{\pi}-\mathcal{E}_{y}(h)]\Big)-\langle f|f\rangle_{\pi}$
$=t\cdot\mathrm{var}(Q_{x},f)+(1-t)\cdot\mathrm{var}(Q_{y},f),$
where the first inequality is strict if (d) holds. However, if (b) and (c) hold, then also (A3) holds and
$\mathrm{var}(Q_{tx+(1-t)y},f)$
$\leq 2\Big(t[2\langle f|h_{tx+(1-t)y}\rangle_{\pi}-\mathcal{E}_{x}(h_{tx+(1-t)y})]+(1-t)[2\langle f|h_{tx+(1-t)y}\rangle_{\pi}-\mathcal{E}_{y}(h_{tx+(1-t)y})]\Big)-\langle f|f\rangle_{\pi}$
$<2\Big(t\sup_{g\in L_{0}^{2}(\pi)}[2\langle f|g\rangle_{\pi}-\mathcal{E}_{x}(g)]+(1-t)\sup_{h\in L_{0}^{2}(\pi)}[2\langle f|h\rangle_{\pi}-\mathcal{E}_{y}(h)]\Big)-\langle f|f\rangle_{\pi}$
$=t\cdot\mathrm{var}(Q_{x},f)+(1-t)\cdot\mathrm{var}(Q_{y},f),$
where the strict inequality follows from Lemma 62. ∎